### v1

Abstract
We extend a previous result [Sutter et al, J. Phys. A: Math. Theor. 57, 475202 (2024)] to give an explicit form of the set of $v$-representable densities on the one-dimensional torus with any fixed number of particles in contact with a heat bath at finite temperature. The particle interaction has to satisfy some mild assumptions but is kept entirely general otherwise. For densities, we consider the Sobolev space $H^{1}$ and exploit the convexity of the functionals. This leads to a broader set of potentials than the usual $L^{p}$ spaces and encompasses distributions. By including temperature and thus considering all excited states in the Gibbs ensemble, Gâteaux differentiability of the thermal universal functional is guaranteed. This yields $v$-representability and it is demonstrated that the given set of $v$-representable densities is even maximal.
I Introduction
Density-functional theory (DFT) is based on the one-to-one relation between densities and potentials. The famous work of Hohenberg and Kohn [1] introduced an energy functional in terms of the density which as such is only well-defined for densities coming from a ground state, a property called ‘pure-state $v$-representability’. The lack of knowledge about this property can be circumvented by using the constrained-search functional as introduced by Percus [2] and Levy [3]. This functional is defined for all densities coming from a $N$-particle wave function ($N$-representability). However, for the method of Kohn and Sham [4] it is still crucial to solve the $v$-representability problem. The idea of this method is to reproduce the ground-state density of an interacting Hamiltonian by a non-interacting Hamiltonian with another external potential. Thus, the density needs to be $v$-representable for an interacting system but at the same time also for a non-interacting one. A possible extension is to allow for mixed states. Levy [5] and Lieb [6] showed that there are ensemble ground-state density which can not be realized by a pure ground-state. Hence the set of $v$-representable densities is extended by allowing for ensemble states and not only pure states. In [7], it is shown that each strictly positive density in $H^{1}(\mathbb{T})$ on the one-dimensional torus can be realized by a distributional potential in $H^{-1}(\mathbb{T})$. Shortly after, Carvalho Corso [8, 9] showed that for an odd number of particles on the one-dimensional torus each ground-state density of a Hamiltonian with a potential in $H^{-1}(\mathbb{T})$ is non-zero. This result completes the full characterization of $v$-representable densities on the one-dimensional torus. Apart from that it was shown that under the same condition for the number of particles, the external potential is uniquely determined by the ground-state density and that the constrained-search functional is differentiable at strictly positive densities. Additionally, Carvalho Corso [10] presented these results for non-interacting systems but without any requirement on the number of particles.
Already in 1965 Mermin [11] extended the work of Hohenberg and Kohn [1] to non-zero temperatures. However the same question about $v$-representability appears. Therefore we are interested in extending our recent work [7] from zero temperature to elevated temperatures. First results about this problem were presented by Chayes et al. [12] for infinite lattice systems. Eschrig [13] formally argued that for the grand canonical ensemble on a $3$-dimensional torus the thermal universal functional is differentiable at $v$-representable densities and that these densities are strictly positive. However, he does not give an explicit expression for the set of $v$-representable densities. In Eschrig’s work, the underlying topology is the usually considered Lebesgue space $L^{3}$, yet the domain of the universal functional (the set of physical densities) is not open in $L^{3}$. In particular, arbitrary close to any physical density there is n density which takes negative values. Hence the universal functional is discontinuous at each physical density. This problem is already extensively studied by Lammert [14] for the zero temperature case. While Lammert is able to show differentiability at certain densities, he does not characterize the set of $v$-representable densities, a problem finally solved on the one-dimensional torus in Sutter et al. [7]. In our previous and the present work, the problems with openness and continuity are circumvented by considering the $H^{1}$ topology instead of the $L^{3}$ topology. Then, the set of $v$-representable densities is open and the universal functional is continuous on that set. Hence, we can investigate Gâteaux differentiability. Another important difference to Eschrig’s work is that we allow for a broader class of potentials and we show that this is even necessary for $v$-representability.
In this work we consider the same class of systems as in Sutter et al. [7] but at elevated temperatures. This means that there are a fixed number of particles on the one-dimensional torus and that we allow for a whole class of interactions including $W=0$. Elevated temperature means that besides the Hamiltonian an additional entropic term has to be considered in the variational principle. In particular, this means that instead of the usual internal energy functional we work with the Helmholtz functional. The minimizer of the Helmholtz functional is explicitly known in terms of the Hamiltonian and it is called Gibbs state. As in [15; 16] the entropy term acts as a regularizer and we are not only able to show that each strictly positive density is $v$-representable but we also get (Gâteaux) differentiability of the universal functional for all interactions in the considered class. Additionally, we show that each density coming from a Gibbs state is strictly positive as Carvalho Corso [8, 9] showed for zero temperature and a odd number of particles. Hence we get the full characterization of $v$-representable densities. The reason for these additional results compared to the zero temperature case with even number of particles, is that the minimizer of Helmholtz functional occupies not only the ground state of the Hamiltonian but also all excited states. Degenerate states get equal weights which yields a unique minimizer of the Helmholtz functional and strict convexity of the Helmholtz free energy functional (in terms of potentials). We formulate the main result of this work.
Theorem 1.
Define the following set of densities on the one-dimensional torus $\mathbb{T}$ and the space of (distributional) one-body potentials,
$\mathscr{X}_{>0}$ $=\{\rho\in L^{2}(\mathbb{T})\mid\nabla\rho\in L^{2}(\mathbb{T}),\smallint\rho=N,\forall x\in\mathbb{T}:\rho(x)>0\}$ (1a)
$\mathscr{X}^{*}$ $=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (1b)
Then for every $\rho\in\mathscr{X}_{>0}$ there is an equivalence class $v]\in\mathscr{X}^{*}$ with corresponding potential $V=\sum_{j=1}^{N}v(x_{j})$, such that $\rho$ is the density of the Gibbs state (see Proposition 3) corresponding to the self-adjoint Hamiltonian $H_{v}=H_{0}+V=-\frac{1}{2}\Delta+W+V$, with $W$ any (interaction) operator that fulfills the KLMN condition (see Eq. (3)). On the other hand, each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. Moreover, the thermal universal functional $F^{\beta}_{\mathrm{DM}}(\rho)$ (defined in Eq. (15b)) is differentiable if and only if $\rho\in\mathscr{X}_{>0}$. This implies that every $\rho\in\mathscr{X}_{>0}$ is even uniquely $v$-representable by a $v]\in\mathscr{X}^{*}$.
The ‘unique $v$-representability’ is equivalent to the famous Hohenberg–Kohn theorem [1]. To prove the stated theorem above, we start by introducing the basic setting for a $N$-particle system on a one-dimensional torus and the relevant spaces for the densities and the potentials (Section II). Afterwards, we introduce the Helmholtz functional. We present known result about this functional and about the Gibbs state together with the introduction of the thermal universal functional (Section III). Then we show the relation between $v$-representability and subgradients (Section IV) followed by an application of the preceding results to conclude $v$-representability (Section V). We then prove the converse statement, that each Gibbs state has indeed a density in the $v$-representable set, to get a full characterization of the set of $v$-representable densities (Section VI). Lastly, we show that the universal functional is Gâteaux differentiable (Section VII). We conclude with the proof of Theorem 1 and an outlook (Section VIII).
II One-dimensional torus setting
We consider a system at an elevated temperature $T$ consisting of $N$ spin $\frac{1}{2}$-particles. The particles are on the one-dimensional torus $\mathbb{T}$, meaning that they are confined to the interval $[0,1]$ where the points 0 and $1$ are identified with each other. The underlying Hilbert space is then
$\mathcal{H}=\bigl{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}.$ (2)
As noted, the considered Hamiltonians are of the form $H_{v}=H_{0}+V$ where $H_{0}$ consists of the kinetic energy operator and an interaction term and where $V$ is the potential. We can generally write $H_{v}$ instead of $H_{[v]}$, even though the considered potentials come from equivalence classes, since adding a constant to the potential just shifts the eigenvalues while not affecting the eigenstates. The interaction is kept general but positive and it needs to fulfill the KLMN condition (Reed and Simon [17], Th. X.17) with respect to the kinetic energy operator, i.e. for each interaction $W$ there are $0\leq a<1$ and $b\geq 0$ such that the following holds for each element in the form domain $Q(-\frac{1}{2}\Delta)$
$|\langle\Psi,W\Psi\rangle|\leq a\langle\Psi,-\tfrac{1}{2}\Delta\Psi\rangle+b\langle\Psi,\Psi\rangle.$ (3)
This condition guarantees that the Hamiltonian $H_{0}=-\tfrac{1}{2}\Delta+W$ remains self-adjoint with the same form domain $Q(H_{0})=Q(-\frac{1}{2}\Delta)$ if we add an interaction to the kinetic energy operator (which is self-adjoint). In Sutter et al. [7, Appendix] this condition was shown for interactions in $L^{1}([-1,1])$ and for distributional interactions of the form $\nabla g$ with $g\in L^{2}([-1,1])$. Here and in the following we use the notion of quadratic forms [18, Sec. VIII.6] instead of linear operators. For the kinetic energy operator the form domain is given by $Q(-\frac{1}{2}\Delta)=\bigl{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}$ with the Sobolev space $H^{1}(\mathbb{T})=\{\Psi\in L^{2}(\mathbb{T})\mid\nabla\Psi\in L^{2}(\mathbb{T})\}$ (like Eq. 1a this means a weak derivative). We get by integration by parts
$\frac{1}{2}\langle\Psi,-\Delta\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\Psi,-\Delta_{j}\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\nabla_{j}\Psi,\nabla_{j}\Psi\rangle.$ (4)
Together with the normalization constraint this gives the set of wave functions
$\mathscr{W}=\{\Psi\in(H^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\mid\|\Psi\|_{\mathcal{H}}=1\}.$ (5)
The one-particle density $\rho_{\Psi}$ is obtained by integrating out all but one of the spatial coordinates of a wave function squared as well as summing over all spin components, i.e.
$\rho_{\Psi}(x)=N\sum_{\sigma_{1},\ldots,\sigma_{N}}\int\lvert\Psi(x\sigma_{1},x_{2}\sigma_{2},\dotsc,x_{N}\sigma_{N})\rvert^{2}\mathrm{d}x_{2}\dotsb\mathrm{d}x_{N}.$ (6)
In the following theorem we collect results from [6] and [7].
Theorem 2.
We define the following set of densities
$\mathscr{I}=\{\rho\in L^{1}(\mathbb{T})\mid\nabla\sqrt{\rho}\in L^{2}(\mathbb{T}),\rho\geq 0,\smallint\rho=N\}.$ (7)
Then, the following holds.
  1. (i)
    Let $\Psi\in\mathscr{W}$. Then $\rho_{\Psi}\in\mathscr{I}$ and it holds
    $\int\left|\nabla\sqrt{\rho_{\Psi}(x)}\right|^{2}\mathrm{d}x\leq 2T(\Psi).$ (8)
  2. (ii)
    There exist two constants $C_{1},C_{2}>0$ such that for all $\rho\in\mathscr{I}$ there is a $\Psi\in\mathscr{W}$ that has $\rho_{\Psi}=\rho$ and the following estimate holds
    $T(\Psi)\leq C_{1}+C_{2}\int\left|\nabla\sqrt{\rho(x)}\right|^{2}\mathrm{d}x.$ (9)
  3. (iii)
    $\mathscr{X}_{>0}\subset\mathscr{I}\subset\mathscr{X}\subset H^{1}(\mathbb{T})\hookrightarrow C^{0}(\mathbb{T})\hookrightarrow L^{\infty}(\mathbb{T})\hookrightarrow L^{3}(\mathbb{T})\hookrightarrow L^{2}(\mathbb{T})\hookrightarrow L^{1}(\mathbb{T})$, where $\mathscr{X}=\{\rho\in H^{1}(\mathbb{T})\mid\smallint\rho=N\}$ and where $X\hookrightarrow Y$ means that $X$ is continuously embedded in $Y$.
Lieb [6] showed that the set $\mathscr{I}$ is convex. This implies that (finite) convex combinations of elements in $\mathscr{I}$ lie again in $\mathscr{I}$. Motivated by the observations for the zero temperature case [7] we identify the space of external potential by $\mathscr{X}^{*}$, the topological dual of $\mathscr{X}$. This space contains equivalence classes, where the elements are from the space $H^{-1}(\mathbb{T})$ and where two elements are in the same class if they differ only by a constant, i.e., like in Eq. (1b),
$\mathscr{X}^{*}=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (10)
The dual pairing of an element $v=f+\nabla g\in H^{-1}(\mathbb{T})$ with an element $\varphi\in H^{1}(\mathbb{T})$ is given by $v(\varphi)=\langle v,\varphi\rangle=\langle f,\varphi\rangle-\langle g,\nabla\varphi\rangle$. Using the dual space of $\mathscr{X}$ for the external potentials guarantees that the potential energy is finite. That Hamiltonians of the form $H=H_{0}+V$ with $V$ defined by a $v\in\mathscr{X}^{*}$ are connected to a self-adjoint operator is shown in [7] by using the KLNM theorem (Reed and Simon [17], Th. X.17).
III Helmholtz free energy and Gibbs state
Considering a system at fixed temperature $T>0$ (through coupling to a heat bath) means that we cannot rely on a variational principle for the internal energy $E_{v}(\Gamma)=\operatorname{Tr}\{\Gamma H_{v}\}$ any more [19]. Instead, we have to consider the Helmholtz free energy
$\Omega_{v}^{\beta}(\Gamma)=E_{v}(\Gamma)-TS(\Gamma)=\operatorname{Tr}\bigl{\{}\Gamma\bigl{(}H_{v}+\beta^{-1}\log(\Gamma)\bigr{)}\bigr{\}},$ (11)
where $S(\Gamma)=-\operatorname{Tr}\{\Gamma\log(\Gamma)\}$ is the entropy and $\beta=T^{-1}$ is the inverse temperature. The new variational principle then yields the Helmholtz functional
$\Omega^{\beta}(v)=\inf_{\Gamma}\Omega^{\beta}_{v}(\Gamma)=\inf_{\Gamma}\operatorname{Tr}\bigl{\{}\Gamma\bigl{(}H_{v}+\beta^{-1}\log(\Gamma)\bigr{)}\bigr{\}}$ (12)
and can been equivalentely seen as the Legendre transform of $E_{v}$ (in terms of entropy) to $\Omega^{\beta}(v)$ (in terms of temperature). Here the infimum is over all density matrices, i.e. over all (possibly infinite) convex sums $\Gamma=\sum_{k}\lambda_{k}\langle\Psi_{k},\cdot\rangle\Psi_{k}$, $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$, with $\Psi_{k}\in\mathscr{W}$ all orthogonal. Making the Helmholtz functional stationary with respect to small perturbation $\delta\Gamma$ with $\operatorname{Tr}\{\delta\Gamma\}=0$ gives the minimizer.
Proposition 3.
The unique minimizer of the Helmholtz functional is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ where $Z(v)=\operatorname{Tr}\{\mathrm{e}^{-\beta H_{v}}\}$ is called the partition function. The state $\Gamma_{v}$ is called the Gibbs state and it does only exist if $Z(v)$ is finite. Moreover, for each $v\in H^{-1}(\mathbb{T})$ the Gibbs state exists and the Helmholtz free energy is given by
$\Omega^{\beta}(v)=-\beta^{-1}\log\bigl{(}Z(v)\bigr{)}.$ (13)
The proof of Proposition 3 can be found in Appendix A. The reason for $Z(v)$ to be finite is that the particles are trapped in a finite volume. The defining parameters of canonical (quantum) thermodynamics are $(N,V,T)$ and all of these parameters need to be finite. For a counterexample [16, Sec. 2.3] consider the hydrogen atom in $\mathbb{R}^{3}$. The energies are $1/(2n^{2})$ with a degeneracy of $n^{2}$ and with $n\in\mathbb{N}$. Then the partition function is
$Z(v)=\sum\limits_{n=1}^{\infty}n^{2}\mathrm{e}^{\beta/(2n^{2})}\geq\sum\limits_{n=1}^{\infty}n^{2}=\infty.$ (14)
Proposition 3 implies that the density of the Gibbs state is an infinite convex combination of elements in $\mathscr{I}$. Hence we cannot directly conclude that its density is contained in $\mathscr{I}$, but we will show this statement with the following proposition.
Proposition 4.
The density of the Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$ for all $v\in H^{-1}$.
The proposition is proven in Appendix B. It tells us that $v$-representable densities at elevated temperatures are contained in $\mathscr{I}$. Hence we can split the minimization of the Helmholtz functional $\Omega^{\beta}_{v}$ as in the zero temperature case into first a variation over densities in $\mathscr{I}$ followed by a minimization over all density matrices which yield such a density,
$\Omega^{\beta}(v)=\inf_{\rho\in\mathscr{I}}\{F^{\beta}_{\mathrm{DM}}(\rho)+\langle v,\rho\rangle\}\quad\text{where}$ (15a)
$F^{\beta}_{\mathrm{DM}}(\rho)=\inf_{\Gamma\mapsto\rho}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}.$ (15b)
The functional $F_{\mathrm{DM}}^{\beta}(\rho)$ is called the universal functional and we can extend it to $\mathscr{X}$ by setting $F_{\mathrm{DM}}^{\beta}(\rho)=\infty$ for all $\rho\in\mathscr{X}\setminus\mathscr{I}$. Note that because of the existence of the Gibbs state, the infimum in the variation over the densities is attained whereas for the second infimum (in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$) we only know about the existence of a minimizer for $v$-representable densities which are characterized in Theorem 1.
As for zero temperature ([6; 7]), we would like to have a convex formulation of thermal density-functional theory to exploit the properties of convex functions. Due to the concavity of the entropy [16, Th. 12] and linearity of $\Gamma\to\rho$ the following proposition follows.
Proposition 5.
The following two statements hold.
  1. (i)
    The Helmholtz functional $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)$ is convex.
  2. (ii)
    The universal functional $\rho\mapsto F^{\beta}_{\mathrm{DM}}(\rho)$ is convex on $\mathscr{X}$.
IV Relation between $v$-representability and subdifferentials
The relation between $v$-representability and non-empty subdifferential was presented in the literature [20; 21; 22]. This relation, together with some additional results from convex analysis, are important ingredients to show our main result Theorem 1. Therefore, we will repeat the relation and the needed results in this section.
For any proper ($f\neq\infty$) convex function $f:\mathscr{X}\to(-\infty,\infty]$ its subdifferential at a point $\rho\in\mathscr{X}$ is defined as the set
$\partial f(\rho)=\{u\in\mathscr{X}^{*}\mid f(\rho)-f(\rho^{\prime})\leq\langle u,\rho-\rho^{\prime}\rangle\;\forall\rho^{\prime}\in\mathscr{X}\}.$ (16)
Note, that if the inequality $f(\rho)-f(\rho^{\prime})\leq\langle u,\rho-\rho^{\prime}\rangle$ cannot be satisfied for any $u\in\mathscr{X}^{*}$ and all $\rho^{\prime}\in\mathscr{X}$ then the subdifferential is empty. An element of $\partial f(\rho)$ is called a subgradient. Moreover, the function $f$ is Gâteaux differentiable at a point $\rho$ if the function $f$ is continuous at $\rho$ and if the subdifferential $\partial f(\rho)$ contains only a single element [23, Prop. 2.40]. In this case, this element is equal to the derivative. If the general convex function is replaced by the universal functional (which is a proper convex functional by Proposition 5), we get for an element $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ and for all $\rho^{\prime}\in\mathscr{X}$ the following inequality
$F_{\mathrm{DM}}^{\beta}(\rho)+\langle v,\rho\rangle\leq F_{\mathrm{DM}}^{\beta}(\rho^{\prime})+\langle v,\rho^{\prime}\rangle.$ (17)
In particular, $\rho$ is a minimizer for the variation in Eq. (15a). Note that this is only a necessary condition for $v$-representability and not a sufficient one. The reason is that for $v$-representability of a density $\rho$ we additionally require that the variation over $\Gamma\mapsto\rho$ in Eq. (15b) has a minimum too, otherwise there is no Gibbs state behind the density. In this section, we show that the subdifferential of $F^{\beta}_{\mathrm{DM}}(\rho)$ at any density $\rho\in\mathscr{X}_{>0}$ is non-empty. In the next section we then show that these densities indeed come from Gibbs states. The two result together give then $v$-representability for all densities in $\mathscr{X}_{>0}$.
To show the existence of a subgradient for a density $\rho\in\mathscr{X}_{>0}$ we proceed as in Sutter et al. [7]. The only result that significantly changes compared to this reference is the local boundedness of the universal functional $F_{\mathrm{DM}}^{\beta}$.
Lemma 6.
There exist three constants $C_{1},C_{2},C_{3}>0$ such that for all $\rho\in\mathscr{I}$ it holds
$C_{1}\leq F_{\mathrm{DM}}^{\beta}(\rho)\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2}.$ (18)
Proof.
For the lower bound of $F_{\mathrm{DM}}^{\beta}(\rho)$ we have
$F_{\mathrm{DM}}^{\beta}(\rho)\geq\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}=\Omega^{\beta}(0)$ (19)
which is greater than -$\infty$ by Proposition 3. For the upper bound we note that the entropic term $\beta S(\Gamma)$ is negative. Hence we get the following inequality
$F_{\mathrm{DM}}^{\beta}(\rho)\leq\inf_{\Gamma\to\rho}\operatorname{Tr}\{\Gamma H_{0}\}\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2}$ (20)
where we used a previous result [7, Lem. 10] for the second inequality. ∎
Theorem 7.
For any $\rho\in\mathscr{X}_{>0}$ the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous and the subdifferential $\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$ is non-empty.
The proof is the same as in Sutter et al. [7, Th. 15].
V Existence of the Minimum
As discussed in the previous section, we do not only need a non-empty subdifferential of $F_{\mathrm{DM}}^{\beta}(\rho)$ to get $v$-representability but also that the minimum in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ (Eq. (15b)) exists. This minimizer must then be the Gibbs state $\Gamma_{v}$, where $v$ is just the potential that yields $\rho$. In the following we show this for all $\rho\in\mathscr{X}_{>0}$. To do so, we make use of the relative entropy which is defined as
$S(A|B)=\operatorname{Tr}\{A(\log A-\log B)+B-A\}$ (21)
for two non-negative trace class operators $A$, $B$ with $\ker B\subset\ker A$. An important property of the relative entropy is lower semi-continuity.
Lemma 8 (Lami and Shirokov [24], Lemma 4).
The relative entropy $(A,B)\mapsto S(A|B)$ is jointly lower semi-continuous with respect to the weak-* topology. (for non-negative trace class operators).
The next lemma relates the Helmholtz functional and the relative entropy. Note that originally we only defined the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)$ for (normalized) density matrices. However, in the same manner it can be defined for any non-negative trace class operator. Recall that $\Gamma_{v}$ is the Gibbs state corresponding to a potential $v$ and it is given by $\mathrm{e}^{-\beta H_{v}}/Z(v)$.
Lemma 9 (Giesbertz and Ruggenthaler [16]).
$S(\Gamma|\Gamma_{v})+\operatorname{Tr}\{\Gamma-\Gamma_{v}\}=\beta\Omega^{\beta}_{v}(\Gamma)-\beta\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}$.
Since the left hand side in Lemma 9 consists of two weak-* lower semi-continuous functions (norm is decreasing under weak-* limits), we get that the Helmholtz functional together with a term proportional to $\operatorname{Tr}\{\cdot\}$ is lower semi-continuous with respect to the weak-* topology. In particular, we get the following lemma.
Lemma 10.
For $\xi>1$ the mapping $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous if $\Omega^{\beta}_{v}(\Gamma_{v})\leq(\beta(\xi-1))^{-1}$. For $\xi<1$ the function is weak-* lower semi-continuous if $\Omega^{\beta}_{v}(\Gamma_{v})\geq(\beta(\xi-1))^{-1}$.
Proof.
From Lemma 9 we get
$\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}=\frac{1}{\beta}S(\Gamma|\Gamma_{v})-\frac{1}{\beta}\operatorname{Tr}\{\Gamma_{v}\}+\operatorname{Tr}\{\Gamma\}\left(\frac{1}{\beta}+(1-\xi)\Omega^{\beta}_{v}(\Gamma_{v})\right).$ (22)
By Lemma 8 we know that the first term $\beta^{-1}S(\Gamma|\Gamma_{v})$ is weak-* lower semi-continuous. The second term $\beta^{-1}\operatorname{Tr}\{\Gamma_{v}\}$ is just a constant. The last term is weak-* lower semi-continuous if $\beta^{-1}+(1-\xi)\Omega^{\beta}_{v}(\Gamma_{v})\geq 0$. Solving for $\Omega^{\beta}_{v}(\Gamma_{v})$ proves the lemma. ∎
Theorem 11.
For any $\rho\in\mathscr{X}_{>0}$ there is a (normalized) density matrix $\Gamma\mapsto\rho$ such that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log\Gamma)\}$. This shows the existence of a minimizer in Eq. (15b).
Proof.
For a density $\rho\in\mathscr{X}_{>0}$ we pick $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ which we know to exists by Theorem 7. By definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ there is a sequence of density matrices $\Gamma_{n}$ each yielding the density $\rho$ and such that $\operatorname{Tr}\{\Gamma_{n}(H_{0}+\beta^{-1}\log\Gamma_{n})\}\to F^{\beta}_{\mathrm{DM}}(\rho)$. Since all $\Gamma_{n}$ are contained in a ball of radius $1$ we get by the Banach–Alaoglu theorem [18, Th. IV.21] that there is a subsequence (in the following also denoted by $\Gamma_{n}$) and a trace class operator $\Gamma_{0}$ such that $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$. Moreover, we know that $\Gamma_{0}$ is non-negative and that $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$. To prove the theorem, we show that $\Gamma_{0}$ is a minimizer. That means that $\Gamma_{0}$ is normalized to $1$, that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ and that it yields the density $\rho$. We start by showing $\operatorname{Tr}\{\Gamma_{0}\}=1$.
For simplicity, we call the minimum of the Helmholtz functional $\Omega^{\beta}=\Omega^{\beta}_{v}(\Gamma_{v})=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$. Because $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$ and $\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}\}=1$ we have
$\Omega^{\beta}-\xi\Omega^{\beta}$ $=F^{\beta}_{\mathrm{DM}}(\rho)+\langle v,\rho\rangle-\xi\Omega^{\beta}=\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}(H_{v}+\beta^{-1}\log\Gamma_{n})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{n}\}$ (23)
$\geq\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\}$
where we have used the fact that the function $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous due to Lemma 10. For the condition of the lemma to be fulfilled we set $\xi=(\beta\Omega^{\beta})^{-1}+1$. This expression is well-defined, since we can choose a representative of $v]\in\mathscr{X}^{*}$ for which $\Omega^{\beta}_{v}(\Gamma_{v})$ is non-zero. So the condition of Lemma 10 is satisfied. Let $c$ be the constant such that $\operatorname{Tr}\{c\Gamma_{0}\}=1$. (Note that the constant $c$ only exist if $\Gamma_{0}\neq 0$. If $\Gamma_{0}=0$ then the right hand-side of Eq. (23) is 0 where the left hand-side is $\beta^{-1}$. Hence $\Gamma_{0}=0$ is not possible.) Since $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$ it holds $c\geq 1$. Then we can compute
$\Omega^{\beta}-\xi\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})-\xi\Omega^{\beta}\operatorname{Tr}\{c\Gamma_{0}\}$ $=\operatorname{Tr}\{c\Gamma_{0}(H_{v}+\beta^{-1}\log(c\Gamma_{0}))\}-\xi\Omega^{\beta}\operatorname{Tr}\{c\Gamma_{0}\}$ (24)
$=c\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}+\beta^{-1}\operatorname{Tr}\{c\Gamma_{0}\}\log c-c\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\}$
$=c(\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\})+\beta^{-1}\log c$
$\leq c(\Omega^{\beta}-\xi\Omega^{\beta})+\beta^{-1}\log c$
where we used the variational principle for $\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})$, the fact that $\operatorname{Tr}\{c\Gamma_{0}\}=1$, and Eq. (23). This inequality can be transformed to
$c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}\geq 0.$ (25)
It is clear that we have equality if $c$ is equal to $1$. The function $f(c)=c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}$ defined on $1,\infty)$ has $f(1)=0$ and the derivative
$f^{\prime}(c)=\Omega^{\beta}-\xi\Omega^{\beta}+\frac{\beta^{-1}}{c}=-\beta^{-1}+\frac{\beta^{-1}}{c}\begin{cases}=0\quad&c=1\\ <0\quad&c>1,\end{cases}$ (26)
where we used $\xi=(\beta\Omega^{\beta})^{-1}+1$. This means that $f(c)$ is strictly decreasing on $1,\infty)$ with a unique maximum at $c=1$. Hence the inequality (25) can only be satisfied if $c=1$ and we conclude that $\operatorname{Tr}\{\Gamma_{0}\}=1$.
Now that we know that the weak-* limit $\Gamma_{0}$ has trace norm equal to $1$ we can show that $\Gamma_{0}\to\rho$. For this we use the same idea as Lieb [6] and show that for any $f\in L^{\infty}(\mathbb{T})$ we have $\operatorname{Tr}\{\Gamma_{n}M_{f}\}\to\operatorname{Tr}\{\Gamma_{0}M_{f}\}$ where $M_{f}$ denotes pointwise multiplication by $f$. This means that the densities of $\Gamma_{n}$ converge weakly to the density of $\Gamma_{0}$ (in $L^{1}$) and by the uniqueness of weak limits we get $\Gamma_{0}\to\rho$. To do this, let $|\sum_{j}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of $\Gamma_{0}$. Since $\operatorname{Tr}\{\Gamma_{0}\}=1$ there is for any $\epsilon>0$ an $N_{\epsilon}\in\mathbb{N}$ such that
$\sum\limits_{j=1}^{N_{\epsilon}}\lambda_{j}>1-\epsilon.$ (27)
We define the projection
$|.P=\sum\limits_{j=1}^{N_{\epsilon}}\ket{\psi_{j}}\bra{\psi_{j}}$. (28)
The operator $P$ is compact and since $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$ in trace class, there exist $N_{P}$ such that $\lvert\operatorname{Tr}\{\Gamma_{n}P\}-\operatorname{Tr}\{\Gamma_{0}P\}\rvert<\epsilon$ for all $n>N_{P}$. It follows that for $n>N_{P}$ we have $\operatorname{Tr}\{\Gamma_{n}P\}>\operatorname{Tr}\{\Gamma_{0}P\}-\epsilon>1-2\epsilon$ and thus $\operatorname{Tr}\{\Gamma_{n}(1-P)\}<2\epsilon$. Then we get that $\lvert\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)\}\rvert<2\epsilon$ for large enough $n$. We can write
$\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}=\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})PM_{f}\}+\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)M_{f}\}.$ (29)
The first term goes to 0 since $PM_{f}$ is a compact operator. The second term is contained in $C\epsilon,C\epsilon]$ for $n$ large enough where the constant $C$ depends only on $f$. Since $\epsilon$ can be arbitrarily small we conclude that $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}\to 0$ and hence $\Gamma_{0}\to\rho$. By Eq. (23) we get $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ which concludes the proof. ∎
Remark 1.
Note that we never use $\rho>0$ and that we can also replace the condition $\rho\in\mathscr{X}_{>0}$ by the condition that the subdifferential of $F_{\mathrm{DM}}^{\beta}$ at $\rho$ is non-empty.
If we combine Theorem 7 and Theorem 11, we get that each density $\rho\in\mathscr{X}_{>0}$ is $v$-representable, which is the first part of Theorem 1. The full proof for Theorem 1 can be found in Section VIII.
VI Closing the Circle
So far we showed that each density in $\mathscr{X}_{>0}$ is $v$-representable. In this section we show that each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. At elevated temperatures, all excited state of the Hamiltonian contribute to the Gibbs state (and hence also to the density) and since these state form an orthonormal basis, it is not possible that the density vanishes somewhere. We formulate this result in the following theorem that closes the circle from strictly positive densities in $\mathscr{X}_{>0}$ to representing potentials in $\mathscr{X}^{*}$ and back to a strictly positive density of the corresponding Gibbs state.
Theorem 12.
The density of any Gibbs state is everywhere non-zero.
Proof.
Let $v\in H^{-1}(\mathbb{T})$. The Gibbs state is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ which can be written as
$|,\Gamma_{v}=\frac{1}{Z(v)}\sum_{j}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$, (30)
where the $\lambda_{j}$ are the eigenvalues of $H_{v}$ with corresponding eigenstates $\psi_{j}$. We denote the density of $\Gamma_{v}$ by $\rho$ and the densities of $\psi_{j}$ by $\rho_{j}$. Let us assume that the density $\rho$ is equal to 0 at $x_{0}$. By the relation
$\rho(x)=\frac{1}{Z(v)}\sum_{j}\mathrm{e}^{-\beta\lambda_{j}}\rho_{j}(x)$ (31)
we conclude that $\rho_{j}(x_{0})=0$ for all $j$ since the weights are all non-zero. This implies that all $\psi_{j}$ vanish on $U=\{x_{0}\}\times[0,1]^{N-1}$. (The restriction of $\psi_{j}$ to $U$ is well-defined by the trace operator for Sobolev spaces [25, p. 315].) Consequently, each linear combination of the $\psi_{j}$ vanishes on $U$. Yet, this statement makes only sense if the linear combination is in $H^{1}$ too. Since the $\psi_{j}$ form a basis of $L^{2}(\mathbb{T}^{N})$ we also need to be able to express each function in $H^{1}(\mathbb{T}^{N})\subset L^{2}(\mathbb{T}^{N})$ as their linear combination. Now this yields a contradiction since clearly not every function in $H^{1}(\mathbb{T}^{N})$ vanishes on U. We note that while the restriction to $U$ is well-defined for $H^{1}(\mathbb{T}^{N})$ functions, this is not the case for functions in $L^{2}(\mathbb{T}^{N})$. ∎
VII Gâteaux differentiability
The last part of the main result (Theorem 1) is about Gâteaux differentiability of $F_{\mathrm{DM}}^{\beta}$ at densities $\rho\in\mathscr{X}_{>0}$. As mentioned in Section IV, Gâteaux differentiability can be achieved by continuity and a single-element subdifferential. Thus, the goal of this section is to show that each density $\rho\in\mathscr{X}_{>0}$ has only one subgradient. We do that by using standard results from convex analysis and by exploiting the following lemma.
Lemma 13.
The Helmholtz free energy $\Omega^{\beta}(v)$ is a strictly concave function in $v\in H^{-1}(\mathbb{T})$.
This result is obtained by exploiting the uniqueness of the Gibbs state. The proof proceeds analogously to [16, Th. 10]. Additionally, we need the following definition and lemmata.
Definition 14 (Barbu and Precupanu [23], p. 71).
Let $f$ be a convex function on $\mathscr{X}$. The closure of $f$, denoted by $\mathrm{cl}f$, is the lower semi-continuous hull of $f$, i.e.  $\mathrm{cl}f(\rho)=\liminf_{\rho^{\prime}\to\rho}f(\rho^{\prime})$ for all $\rho\in\mathscr{X}$ if $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})>$-$\infty$ for all $\tilde{\rho}$ and it is the constant function -$\infty$ if there is a $\tilde{\rho}\in\mathscr{X}$ such that $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})=$-$\infty$.
Since the universal functional $F_{\mathrm{DM}}^{\beta}$ is uniformly bounded from below by Lemma 6, the closure of it is given by $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$. Another characterization of the closure of a convex function is by its biconjugate as the following lemma shows.
Lemma 15 (Barbu and Precupanu [23], Cor. 2.23).
The biconjugate of a convex function $f$ coincides with its closure. That is, $f^{**}=\mathrm{cl}f$.
Lemma 16 (Barbu and Precupanu [23], Prop. 2.33).
Let $f:\mathscr{X}\to\mathbb{R}\cup\{\infty\}$ be a proper, convex and lower semi-continuous function and $f^{*}$ its conjugate. Then the following two statements are equivalent,
  1. (i)
    $v\in\partial f(\rho)$
  2. (ii)
    $\rho\in\partial f^{*}(v)$.
Lemma 17 (Barbu and Precupanu [23], Prop. 2.40).
If a convex function $f$ on $\mathscr{X}$ is Gâteaux differentiable at $\rho$, then $\partial f(\rho)$ consists of a single element $v=\mathrm{grad}f(\rho)$. Conversely, if $f$ is continuous at $\rho$ and if $\partial f(\rho)$ contains only a single element, then $f$ is Gâteaux differentiable at $\rho$ and $\mathrm{grad}f(\rho)=\partial f(\rho)$.
We have now all the needed ingredients to show Gâteaux differentiability of the universal functional.
Theorem 18.
The universal functional $F_{\mathrm{DM}}^{\beta}$ is Gâteaux differentiable at any density $\rho\in\mathscr{X}_{>0}$.
Proof.
According to Theorem 7, the functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$. By Lemma 17 it is then enough to show that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued for all densities $\rho\in\mathscr{X}_{>0}$. The conjugate function of $F_{\mathrm{DM}}^{\beta}$ is given by $\Omega^{\beta}(-\cdot)$. If we conjugate this function we end up with the biconjugate of $F_{\mathrm{DM}}^{\beta}$ which is defined by
$F_{\mathrm{DM}}^{\beta**}(\rho)=\sup\big{\{}\Omega^{\beta}(v)-\langle\rho,v\rangle\mid v\in H^{-1}(\mathbb{T})\big{\}}.$ (32)
The biconjugate $F_{\mathrm{DM}}^{\beta**}$ is a convex, lower semi-continuous function and it coincides with the convex envelope of $F_{\mathrm{DM}}^{\beta}$. Since $F_{\mathrm{DM}}^{\beta}$ is convex, Lemma 15 states that the closure of $F_{\mathrm{DM}}^{\beta}$ equals to its biconjugate $F_{\mathrm{DM}}^{\beta**}$. By Definition 14 and its proceeding discussion $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$. Theorem 7 shows that the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$ and hence we get that for each $\rho>0$
$F_{\mathrm{DM}}^{\beta**}(\rho)=\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})=F_{\mathrm{DM}}^{\beta}(\rho).$ (33)
This means that $F_{\mathrm{DM}}^{\beta}$ and its biconjugate $F_{\mathrm{DM}}^{\beta**}$ coincide on the open set $\mathscr{X}_{>0}$. In particular this means that the subdifferentials are equal for densities $\rho\in\mathscr{X}_{>0}$, i.e. $\partial F_{\mathrm{DM}}^{\beta}(\rho)=\partial F_{\mathrm{DM}}^{\beta**}(\rho)$. So instead of showing that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued, we show it for $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$. Note that the conjugate function of $F_{\mathrm{DM}}^{\beta**}$ is $\Omega^{\beta}(-\cdot)$ since the triple conjugate of $F_{\mathrm{DM}}^{\beta}$ is equal to the conjugate. The function $F_{\mathrm{DM}}^{\beta**}$ fulfills all assumptions of Lemma 16 and hence we get
$v\in\partial F_{\mathrm{DM}}^{\beta**}(\rho)\Longleftrightarrow\rho\in\partial\left(-\Omega^{\beta}\right)(v).$ (34)
By Lemma 13 we have that $\Omega^{\beta}$ is strictly convex. In particular this means that a density $\rho$ can only be contained in one subdifferential. Otherwise, if $\rho$ is contained in the subdifferential at $v_{1}$ and $v_{2}$, then the function $\Omega^{\beta}$ would not be strictly convex on the line with boundary points $v_{1}$, $v_{2}$. By the relation (34) we get that $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$ is single valued. This concludes the proof. ∎
VIII Summary and outlook
We can now collect the results from the previous sections to finally give a proof of our main result.
Proof of Theorem 1.
Let $\rho\in\mathscr{X}_{>0}$, then Theorem 7 shows that $\partial F^{\beta}_{\mathrm{DM}}(\rho)$ is non-empty. This implies that $\rho$ is a minimizer for the variational principle in Eq. (15a), where the negative potential is in the subdifferential, i.e., $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$. The minimizer of Eq. (15b), guaranteed to exist by Theorem 11, is then also a minimizer in the Helmholtz functional (Eq. (12)). This minimizer is unique by Proposition 3 and is equal to the Gibbs state $\Gamma_{v}$. We conclude, that $\rho$ is the density of $\Gamma_{v}$ and hence $v$-representable. On the other hand, every density of a Gibbs state $\Gamma_{v}$ with $v\in H^{-1}(\mathbb{T})$ is contained in $\mathscr{I}$ according to Proposition 4 and is strictly positive by Theorem 12. In particular, every $v$-representable density is in $\mathscr{X}_{>0}$ and $F^{\beta}_{\mathrm{DM}}$ is not differentiable at densities in $\mathscr{I}\setminus\mathscr{X}_{>0}$. On the other hand, Theorem 18 shows that the universal functional $F^{\beta}_{\mathrm{DM}}$ is Gâteaux differentiable at any $\rho\in\mathscr{X}_{>0}$. ∎
In this work we provided an explicit characterization of the set of uniquely $v$-representable densities on the one-dimensional torus at elevated temperatures and established (Gâteaux) differentiability of the universal functional. In the same manner the theory can be developed for any other one-dimensional bounded domain. As an extensive parameter, the volume needs to be finite, and thus a periodic setting is the natural choice for elevated temperatures. The restriction to one dimension remains as the main limitation of the current formulation. Another interesting observation is that we do need distributional potentials. In DFT often the potential set $L^{3/2}+L^{\infty}$ is considered [6]. However, for zero temperature and one particle it was already observed by Englisch and Englisch [26] and Chayes et al. [12] that relatively nice densities need distributional potentials. Sutter et al. [7] showed that distributional potentials are sufficient for $v$-representability of the set $\mathscr{X}_{>0}$. For elevated temperature we further show that distributional potential are also necessary. This leads to the clear conjecture that also for higher dimensional domains some sort of distributional potentials are needed. However, the continuous Sobolev embedding $H^{1}\hookrightarrow L^{\infty}$ is crucial and it holds only for one-dimensional domains. For higher dimension we either need to require higher regularity or higher integrability conditions. This means that for the Sobolev space $W^{m,p}$, we need to have $m$ and $p$ large enough. The problem is then that not all Gibbs state densities might be contained in $W^{m,p}$ and that a Hamiltonian with a potential in $W^{-m,q}=(W^{m,p})^{\ast}$ might not be self-adjoint. Another open question is how to extend the results to the grand canonical ensemble of variable particle number. In this case, the particle-number operator is not bounded anymore and this is currently needed in the proof of Theorem 11 where we show that the weak-* limit yields the same density as the sequence. Additionally, we do not know if the square root of the Gibbs state density is in $H^{1}$. However, we think that the rest of the theory does not fundamentally change and that we get the same conclusion as for the canonical ensemble.
Acknowledgements.
MP acknowledges support from the German Research Foundation under Grant SCHI 1476/1-1 and from ERC-2021-STG grant agreement No. 101041487 REGAL. The authors SMS and KJHG thank The Netherlands Organization for Scientific Research, NWO, for its financial support under Grant No. OCENW.KLEIN.434 and Vici Grant No. 724.017.001. RvL acknowledges the Academy of Finland grant under project number 356906. MR acknowledges the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG), EXC 2056, project ID 390715994.
Appendix A Proof of Proposition 3
We still need to show that Proposition 3 holds. The result that the unique minimizer of the Helmholtz functional is given by the Gibbs state $\Gamma_{v}=\mathrm{e}^{\beta H_{v}}/Z(v)$ and the relation $\Omega^{\beta}_{v}(\Gamma_{v})=-\beta\log Z(v)$ is shown in standard literature. Yet, we give here a different approach than Mermin [11] and a more elaborated version than our previous works [16; 15].
Proof of Proposition 3.
To determine the minimizer of the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)=\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$ we make it stationary with respect to small perturbations $\delta\Gamma$ that obey $\operatorname{Tr}\{\delta\Gamma\}=0$. We compute
$\Omega^{\beta}_{v}(\Gamma+\epsilon\delta\Gamma)-\Omega^{\beta}_{v}(\Gamma)=\epsilon\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{(\Gamma+\epsilon\delta\Gamma)\log(\Gamma+\epsilon\delta\Gamma)\}-\beta^{-1}\operatorname{Tr}\{\Gamma\log\Gamma\}.$ (35)
We use the Taylor expansion of $x\log(x)$ to get $(\Gamma+\epsilon\delta\Gamma)\log(\Gamma+\epsilon\delta\Gamma)=\Gamma\log\Gamma+\log(\Gamma)\epsilon\delta\Gamma+\epsilon\delta\Gamma+\mathcal{O}(\epsilon^{2})$. Dividing by $\epsilon$ turns Eq. (35) into
$\frac{\Omega^{\beta}_{v}(\Gamma+\epsilon\delta\Gamma)-\Omega^{\beta}_{v}(\Gamma)}{\epsilon}$ $=\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\log\Gamma\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\}+\mathcal{O}(\epsilon)$ (36)
$=\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\log\Gamma\}+\mathcal{O}(\epsilon).$
The quotient in Eq. (36) must go to zero for any minimizer $\Gamma_{v}$ as $\epsilon\to 0$. Hence we have $\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\log\Gamma_{v}\}=\operatorname{Tr}\{\delta\Gamma(H_{v}+\beta^{-1}\log\Gamma_{v})\}=0$ for all $\delta\Gamma$ which requires $H_{v}+\beta^{-1}\log\Gamma_{v}=\mathrm{const}$. This then leads to the unique minimizer $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ where $Z(v)$ is determined from the normalization constraint.
To get the expression for the Helmholtz free energy, let us first assume that $Z(v)$ is finite, i.e., that the Gibbs state exists. Using the explicit expression of the Gibbs state $\Gamma_{v}$ and $\log\Gamma_{v}=-\beta H_{v}-\log Z(v)$ gives the expression
$\Omega^{\beta}(v)=\Omega^{\beta}_{v}(\Gamma_{v})=\operatorname{Tr}\{\Gamma_{v}(H_{v}+\beta^{-1}\log\Gamma_{v})\}=\operatorname{Tr}\{\Gamma_{v}(H_{v}-H_{v}-\beta^{-1}\log Z(v))\}=-\beta^{-1}\log Z(v).$ (37)
For the last remaining part of the proposition we need to show that $Z(v)$ is indeed finite for any potential $v\in H^{-1}(\mathbb{T})$. First, we show that the partition function is finite for a non-interacting system with zero potential, i.e., $H_{v}=T$. We denote this partition function by $Z_{0}(0)$. The eigenstates of the kinetic-energy operator $\frac{1}{2}\Delta$ on $\mathbb{T}$ are Slater determinants built from the functions $\mathrm{e}^{\mathrm{i}2\pi p\cdot x}$ with $p\in\mathbb{Z}$ which have eigenvalues $2\pi^{2}p^{2}$. We estimate
$Z_{0}(0)=\operatorname{Tr}\{\mathrm{e}^{-\beta T}\}\leq\sum\limits_{p_{1},\ldots,p_{N}\in\mathbb{Z}}\mathrm{e}^{-\beta 2\pi^{2}(p_{1}^{2}+\ldots+p_{N}^{2})}=\Big{[}\sum\limits_{p\in\mathbb{Z}}\mathrm{e}^{-\beta 2\pi^{2}p^{2}}\Big{]}^{N}<\infty.$ (38)
To show that $Z(v)$ is finite for a general interaction and a general potential we first show that $\Omega^{\beta}(v)$ is finite. For that recall that we assume that the interaction satisfies the KLMN condition (3). Moreover, Sutter et al. [7, Lem. 16] showed that any potential $v\in H^{-1}$ fulfills also the KLMN condition $|V(\Psi,\Psi)|\leq aT(\Psi)+b$ where $a>0$ can be made arbitrary small. Here $V(\cdot,\cdot)$ is the quadratic form produced by $v\in H^{-1}$. Combing these two KLMN conditions guarantees the existence of $0<a<1$ and $b$ such that $\langle\Psi,H_{v}\Psi\rangle\geq(1-a)T(\Psi)-b$ for all $\Psi\in\mathscr{W}$. This condition translates to density matrices and we get
$\Omega^{\beta}(v)=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log(\Gamma))\}\geq(1-a)\inf_{\Gamma}\operatorname{Tr}\{\Gamma(T+\big{(}\beta(1-a)\big{)}^{-1}\log(\Gamma))\}-b=-\beta^{-1}\log\big{(}Z_{0}(0)\big{)}-b>-\infty.$ (39)
Let $|H_{v}=\sum_{j=1}^{\infty}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of the Hamiltonian $H_{v}$ which exist due to the compact embedding of $Q(H_{v})=\big{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$ into $\big{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$. We define a sequence of density matrices by
[…]
and we compute
$\Omega^{\beta}(v)<\operatorname{Tr}\{\Gamma_{n}(H_{v}+\beta^{-1}\log(\Gamma_{n}))=\frac{1}{Z_{n}[v]}\sum\limits_{j=1}^{n}\mathrm{e}^{-\beta\lambda_{j}}(-\beta^{-1}\log Z_{n}[v])=-\beta^{-1}\log Z_{n}[v].$ (41)
The left-hand side is bounded from below by Eq. (39) and we get that $Z_{n}[v]$ is uniformly bounded from above. This implies $Z(v)<\infty$. ∎
Appendix B Proof of Proposition 4
In this section we show that the density of a Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$. To do so, note that as explained in the proof of Proposition 3 we have for each potential $v$ two constants $0\leq a<1$ and $b\geq 0$ such that
$(1-a)T-b\leq H_{v}\leq(1+a)T+b$ (42)
where $T=-1/2\Delta$ is the kinetic energy operator. Moreover, the spectrum of the Hamiltonian $H_{v}$ (and also of $T$) contains only discrete values due to the compact embedding $Q(H_{v})=H^{1}(\mathbb{T}^{N})\hookrightarrow L^{2}(\mathbb{T}^{N})$. Then we can express $H_{v}$ and $T$ in spectral decomposition
[…]
where we additionally require $\lambda_{1}\leq\lambda_{2}\leq\ldots$ and $\mu_{1}\leq\mu_{2}\leq\ldots$. Then, we get the following lemma.
Lemma 19.
The $n$’th eigenvalue of $H_{v}$ has the lower and upper bound, $(1-a)\mu_{n}-b\leq\lambda_{n}\leq(1+a)\mu_{n}+b$.
Proof.
Let $\phi_{j}$ be the states introduced in Eq. (43). By the min-max-principle for self-adjoint operators and by Eq. (42) we get
[…]
For the lower bound we proceed in the same way but we use the max-min-principle instead to get
[…]
We use the lower and upper bound of the eigenvalues to proof that the kinetic energy expectation is finite.
Lemma 20.
For all, $v\in H^{-1}$ we have $\mathrm{Tr}\{T\Gamma_{v}\}<\infty$.
Proof.
The Gibbs state is given by $|\Gamma_{v}=\frac{1}{Z(v)}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$. We first show that the energy expectation is finite by using Lemma 19. That is,
-$\infty$ $<\mathrm{Tr}\{H_{v}\Gamma_{v}\}=\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\lambda_{j}\mathrm{e}^{-\beta\lambda_{j}}\leq\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\big{[}(1+a)\mu_{j}+b\big{]}\mathrm{e}^{-\beta[(1-a)\mu_{j}-b]}$
$=\frac{(1+a)}{Z(v)}\mathrm{e}^{\beta b}\sum\limits_{j=1}^{\infty}\mu_{j}\mathrm{e}^{-\beta(1-a)\mu_{j}}+\frac{\mathrm{e}^{\beta b}b}{Z(v)}\sum\limits_{j=1}^{\infty}\mathrm{e}^{-\beta(1-a)\mu_{j}}.$
A similar computation as in Eq. (38) shows that the two sums are finite and we get $\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty$. Using Eq. (42) and positivity of the kinetic energy operator gives
$-b\leq(1-a)\mathrm{Tr}\{T\Gamma_{v}\}-b\leq\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty,$
which concludes the proof. ∎
We can now prove Proposition 4.
Proof of Proposition 4.
For simplicity we introduce $\alpha_{j}=\mathrm{e}^{-\beta\lambda_{j}}/Z(v)$ such that the Gibbs state is $|.\Gamma_{v}=\sum_{j=1}^{\infty}\alpha_{j}\ket{\psi_{j}}\bra{\psi_{j}}$. Then, its density is given by $\rho=\sum_{j=1}^{\infty}\alpha_{j}\rho_{j}$ with $\psi_{j}\mapsto\rho_{j}$. By Theorem 2 all the pure state densities $\rho_{j}$ are contained in $\mathscr{I}$. We define a sequence $\rho^{n}=\sum_{j=1}^{n}\alpha_{j}\rho_{j}$. Clearly, $\rho^{n}\to\rho$ in $L^{1}(\mathbb{T})$ and $\|\sqrt{\rho^{n}}\rVert_{2}^{2}\leq N$. Since $\int(\nabla\sqrt{\rho})^{2}\mathrm{d}x$ is convex on $\mathscr{I}$ [6, remark after Eq. (1.10)], we find by Jensen’s inequality and by Theorem 2
$\int(\nabla\sqrt{\rho^{n}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}\int(\nabla\sqrt{\rho_{j}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}2T(\psi_{j})\leq 2\mathrm{Tr}\{T\Gamma_{v}\}<\infty,$
where we get a uniform bound by Lemma 20. In particular, $\sqrt{\rho^{n}}$ is a bounded sequence in $H^{1}(\mathbb{T})$ and by the Banach–Alaoglu theorem, there is a subsequence (also denoted by $\sqrt{\rho^{n}}$) and $f\in H^{1}(\mathbb{T})$ such that $\sqrt{\rho^{n}}\rightharpoonup f$ weakly in $H^{1}$. But since $H^{1}(\mathbb{T})$ is compactly embedded in $L^{2}(\mathbb{T})$ there is also a subsequence (again denoted by $\sqrt{\rho^{n}}$) such that $\sqrt{\rho^{n}}\to f$ strongly in $L^{2}(\mathbb{T})$. Then, we have
$\|\rho^{n}-f^{2}\rVert_{1}=\int(\sqrt{\rho^{n}}-f)(\sqrt{\rho^{n}}+f)\mathrm{d}x\leq\|\sqrt{\rho^{n}}-f\rVert_{2}\|\sqrt{\rho^{n}}+f\rVert_{2}\to 0$
since $\|\sqrt{\rho^{n}}+f\rVert_{2}$ is bounded. This means that $\rho^{n}\to f^{2}$ strongly in $L^{1}$ and we conclude $\sqrt{\rho}=f\in H^{1}(\mathbb{T})$. ∎

### v2

@@
 The particle interaction has to satisfy some mild assumptions but is kept entirely general otherwise.
-For densities, we consider the Sobolev space $H^{1}$ and exploit the convexity of the functionals.
-This leads to a broader set of potentials than the usual $L^{p}$ spaces and encompasses distributions.
-By including temperature and thus considering all excited states in the Gibbs ensemble, Gâteaux differentiability of the thermal universal functional is guaranteed.
-This yields $v$-representability and it is demonstrated that the given set of $v$-representable densities is even maximal.
+Since we find the necessary domain for densities to be the Sobolev space $H^{1}$, this leads to a broader dual space for potentials than the usual $L^{p}$ spaces and encompasses distributions.
+By including temperature and thus considering all excited states in the Gibbs ensemble, Gâteaux differentiability of the thermal universal functional is achieved.
+This is equivalent to unique $v$-representability and we can prove that the given set of $v$-representable densities is even maximal.
 I Introduction
 Density-functional theory (DFT) is based on the one-to-one relation between densities and potentials.
-The famous work of Hohenberg and Kohn [1] introduced an energy functional in terms of the density which as such is only well-defined for densities coming from a ground state, a property called ‘pure-state $v$-representability’.
-The lack of knowledge about this property can be circumvented by using the constrained-search functional as introduced by Percus [2] and Levy [3].
-This functional is defined for all densities coming from a $N$-particle wave function ($N$-representability).
+The famous work of Hohenberg and Kohn [21] introduced an energy functional in terms of the density, which as such is only well-defined for densities coming from a ground state, a property called ‘pure-state $v$-representability’.
+The lack of knowledge about this property, termed the “$v$-representability problem”, can be circumvented by using the constrained-search functional as introduced by Percus [33] and Levy [26].
+This functional is defined for all densities coming from an $N$-particle wave function with finite kinetic energy ($N$-representability).
 However, for the method of Kohn and Sham [23] it is still crucial to solve the $v$-representability problem.
 The idea of this method is to reproduce the ground-state density of an interacting Hamiltonian by a non-interacting Hamiltonian with another external potential.
-Thus, the density needs to be $v$-representable for an interacting system but at the same time also for a non-interacting one.
+Thus, the density needs to be simultaneously $v$-representable for an interacting and a non-interacting system.
 A possible extension is to allow for mixed states.
-Levy [5] and Lieb [6] showed that there are ensemble ground-state density which can not be realized by a pure ground-state.
+Levy [27] and Lieb [28] showed that there are ensemble ground-state densities which cannot be realized by a pure ground state.
 Hence the set of $v$-representable densities is extended by allowing for ensemble states and not only pure states.
-In [7], it is shown that each strictly positive density in $H^{1}(\mathbb{T})$ on the one-dimensional torus can be realized by a distributional potential in $H^{-1}(\mathbb{T})$.
+In Sutter et al.
+[38], it is shown that each strictly positive density in $H^{1}(\mathbb{T})$ on the one-dimensional torus can be realized by a distributional potential in $H^{-1}(\mathbb{T})$.
 Shortly after, Carvalho Corso [7, 8] showed that for an odd number of particles on the one-dimensional torus each ground-state density of a Hamiltonian with a potential in $H^{-1}(\mathbb{T})$ is non-zero.
-This result completes the full characterization of $v$-representable densities on the one-dimensional torus.
-Apart from that it was shown that under the same condition for the number of particles, the external potential is uniquely determined by the ground-state density and that the constrained-search functional is differentiable at strictly positive densities.
-Additionally, Carvalho Corso [10] presented these results for non-interacting systems but without any requirement on the number of particles.
+This result completes the full characterization of $v$-representable densities in this case.
+The same references also show that the external potential is uniquely determined by the ground-state density and that the constrained-search functional is differentiable at strictly positive densities.
+Additionally, Carvalho Corso [9] presents these results for non-interacting systems but without any requirement on the number of particles.
 Already in 1965 Mermin [30] extended the work of Hohenberg and Kohn [21] to non-zero temperatures.
-However the same question about $v$-representability appears.
-Therefore we are interested in extending our recent work [7] from zero temperature to elevated temperatures.
-First results about this problem were presented by Chayes et al.
+Recently, there has been renewed, broad interest in the field of thermal DFT, ranging from the application of DFT to warm dense matter [22; 36; 18] and Hubbard models [31] to more theoretical work concerning exact conditions [5], the adiabatic connection [19], a rigorous proof of the Hohenberg–Kohn theorem for systems with temperature [16], and the formulation of Entropic DFT [39].
+For this reason, we want to try and answer the same question about $v$-representability for the case of thermal ensembles and hereby extend our recent work [38] from zero temperature to elevated temperatures.
+First results about this problem were achieved by Chayes et al.
 [10] for infinite lattice systems.
@@
 However, he does not give an explicit expression for the set of $v$-representable densities.
-In Eschrig’s work, the underlying topology is the usually considered Lebesgue space $L^{3}$, yet the domain of the universal functional (the set of physical densities) is not open in $L^{3}$.
-In particular, arbitrary close to any physical density there is n density which takes negative values.
+In Eschrig’s work, the underlying topology is the usually considered Lebesgue space $L^{3}$, yet the domain of the thermal universal functional (the set of physical densities) is not open in $L^{3}$.
+The same problem persists for the usual universal functional of DFT.
+In particular, arbitrary close to any physical density there is a density that takes negative values.
 Hence the universal functional is discontinuous at each physical density.
-This problem is already extensively studied by Lammert [14] for the zero temperature case.
-While Lammert is able to show differentiability at certain densities, he does not characterize the set of $v$-representable densities, a problem finally solved on the one-dimensional torus in Sutter et al.
-[7].
-In our previous and the present work, the problems with openness and continuity are circumvented by considering the $H^{1}$ topology instead of the $L^{3}$ topology.
-Then, the set of $v$-representable densities is open and the universal functional is continuous on that set.
-Hence, we can investigate Gâteaux differentiability.
-Another important difference to Eschrig’s work is that we allow for a broader class of potentials and we show that this is even necessary for $v$-representability.
-In this work we consider the same class of systems as in Sutter et al.
+This problem was previously studied extensively by Lammert [25] for the zero temperature case.
+While Lammert was able to show differentiability at certain densities by considering a Sobolev space topology, he did not characterize the set of all $v$-representable densities, a problem finally solved in Sutter et al.
+[38] on the one-dimensional torus.
+In our previous and the present work, the problems with openness and continuity are circumvented by considering the finer $H^{1}$ topology instead of the $L^{3}$ topology.
+A topology is finer than another if it has more open sets, or equivalently, less convergent sequences.
+In one dimension it holds in terms of norms that for any $2\leq p\leq\infty$ there is a $C>0$ such that for all $f\in H^{1}$ one has
+$\|f\|_{L^{p}}\leq C\|f\|_{H^{1}}$ (1)
+and one says that $H^{1}$ is continuously embedded into $L^{p}$ [1, Th.
+4.12, Part I.A].
+Consequently, in one dimension the $H^{1}$ topology is finer than all $L^{p}$ topologies, $p\geq 2$.
+This leads to the set of strictly positive densities in $H^{1}$ being open and to a universal functional that is continuous on that set.
+Hence, we can subsequently investigate Gâteaux differentiability, the property that for a functional $f:X\to\mathbb{R}$ on a Banach space $X$ at every point $x\in U$ from an open domain $U\subseteq X$ the functional derivative
+$\nabla f(x)=\lim_{\tau\to 0}\frac{f(x+\tau y)-f(x)}{\tau}$ (2)
+exists for all directions $y\in X$.
+In this work, we study the same class of systems as in Sutter et al.
 [38] but at elevated temperatures.
-This means that there are a fixed number of particles on the one-dimensional torus and that we allow for a whole class of interactions including $W=0$.
-Elevated temperature means that besides the Hamiltonian an additional entropic term has to be considered in the variational principle.
-In particular, this means that instead of the usual internal energy functional we work with the Helmholtz functional.
-The minimizer of the Helmholtz functional is explicitly known in terms of the Hamiltonian and it is called Gibbs state.
-As in [15; 16] the entropy term acts as a regularizer and we are not only able to show that each strictly positive density is $v$-representable but we also get (Gâteaux) differentiability of the universal functional for all interactions in the considered class.
-Additionally, we show that each density coming from a Gibbs state is strictly positive as Carvalho Corso [8, 9] showed for zero temperature and a odd number of particles.
+This means that we consider a fixed number of fermionic particles on the one-dimensional torus and a very general class of interactions, including of course $W=0$.
+Elevated temperature means that besides the Hamiltonian an additional entropic term has to be added in the variational principle, and that instead of the usual energy we work with the Helmholtz free energy.
+The minimizer of the Helmholtz free energy is explicitly known in terms of the Hamiltonian and it is called Gibbs state.
+As demonstrated before [37; 17], the entropy term acts as a regularizer and we are not only able to show that each strictly positive density is $v$-representable but we also get (Gâteaux) differentiability of the thermal universal functional for all interactions in the considered class.
+Additionally, we show that each density coming from a Gibbs state is strictly positive.
+The same is true for zero temperature and an odd number of particles [7; 8].
 Hence we get the full characterization of $v$-representable densities.
-The reason for these additional results compared to the zero temperature case with even number of particles, is that the minimizer of Helmholtz functional occupies not only the ground state of the Hamiltonian but also all excited states.
-Degenerate states get equal weights which yields a unique minimizer of the Helmholtz functional and strict convexity of the Helmholtz free energy functional (in terms of potentials).
-We formulate the main result of this work.
+The reason for these additional results is that the minimizer of the Helmholtz free energy does not only occupy the ground state of the Hamiltonian but partially also all excited states.
+Degenerate states get equal weights and this yields a unique minimizer and strict convexity of the Helmholtz functional (in terms of potentials).
+We now formulate the main result of this work.
 Theorem 1.
-Define the following set of densities on the one-dimensional torus $\mathbb{T}$ and the space of (distributional) one-body potentials,
-$\mathscr{X}_{>0}$ $=\{\rho\in L^{2}(\mathbb{T})\mid\nabla\rho\in L^{2}(\mathbb{T}),\smallint\rho=N,\forall x\in\mathbb{T}:\rho(x)>0\}$ (1a)
-$\mathscr{X}^{*}$ $=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (1b)
-Then for every $\rho\in\mathscr{X}_{>0}$ there is an equivalence class $v]\in\mathscr{X}^{*}$ with corresponding potential $V=\sum_{j=1}^{N}v(x_{j})$, such that $\rho$ is the density of the Gibbs state (see Proposition 3) corresponding to the self-adjoint Hamiltonian $H_{v}=H_{0}+V=-\frac{1}{2}\Delta+W+V$, with $W$ any (interaction) operator that fulfills the KLMN condition (see Eq.
+Define a set of densities and a space of (distributional) one-body potentials on the one-dimensional torus $\mathbb{T}$, $\mathscr{X}_{>0}$ $=\{\rho\in L^{2}(\mathbb{T})\mid\nabla\rho\in L^{2}(\mathbb{T}),\smallint\rho=N,\forall x\in\mathbb{T}:\rho(x)>0\}$ (3a) $\mathscr{X}^{*}$ $=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (3b) Then for every $\rho\in\mathscr{X}_{>0}$ there is a unique potential $v$ with equivalence class $v]\in\mathscr{X}^{*}$, such that $\rho$ is the density of the Gibbs state (see Proposition 3) corresponding to the self-adjoint Hamiltonian $H_{v}=H_{0}+V=-\frac{1}{2}\Delta+W+V,\quad V=\sum_{j=1}^{N}v(x_{j}),$ (4) where $W$ is any (interaction) operator that fulfills the KLMN condition (see Eq.
 (10)).
@@
 Moreover, the thermal universal functional $F^{\beta}_{\mathrm{DM}}(\rho)$ (defined in Eq.
-(15b)) is differentiable if and only if $\rho\in\mathscr{X}_{>0}$.
-This implies that every $\rho\in\mathscr{X}_{>0}$ is even uniquely $v$-representable by a $v]\in\mathscr{X}^{*}$.
-The ‘unique $v$-representability’ is equivalent to the famous Hohenberg–Kohn theorem [1].
-To prove the stated theorem above, we start by introducing the basic setting for a $N$-particle system on a one-dimensional torus and the relevant spaces for the densities and the potentials (Section II).
-Afterwards, we introduce the Helmholtz functional.
-We present known result about this functional and about the Gibbs state together with the introduction of the thermal universal functional (Section III).
-Then we show the relation between $v$-representability and subgradients (Section IV) followed by an application of the preceding results to conclude $v$-representability (Section V).
-We then prove the converse statement, that each Gibbs state has indeed a density in the $v$-representable set, to get a full characterization of the set of $v$-representable densities (Section VI).
-Lastly, we show that the universal functional is Gâteaux differentiable (Section VII).
-We conclude with the proof of Theorem 1 and an outlook (Section VIII).
+(22b)) is Gâteaux differentiable if and only if $\rho\in\mathscr{X}_{>0}$.
+This implies that for a fixed interaction $W$, every $\rho\in\mathscr{X}_{>0}$ is uniquely $v$-representable by a $v]\in\mathscr{X}^{*}$.
+The property of ‘unique $v$-representability’ is equivalent to the famous Hohenberg–Kohn theorem [21; 15].
+To prove the stated theorem above, we start by introducing the basic setting for an $N$-particle system on a one-dimensional torus and the relevant spaces for densities and potentials (Section II).
+Afterwards, we introduce the Helmholtz free energy and present known results about it and its unique minimizer, the Gibbs state, together with the definition of the thermal universal functional (Section III).
+Next, we show the non-emptiness of the subdifferential of the thermal universal functional that is a necessary condition for $v$-representability (Section IV).
+This is followed by an application of the preceding results to conclude $v$-representability (Section V).
+We then prove the converse statement, that each Gibbs state indeed yields a density in $\mathscr{X}_{>0}$, to achieve a full characterization of the set of $v$-representable densities (Section VI).
+Lastly, we show that the thermal universal functional is Gâteaux differentiable (Section VII).
+We conclude with the proof of Theorem 1 that is just a collection of previous results and an outlook (Section VIII).
+Table 1: Notation Table
+Symbol Description
+$\mathbb{T}$ one-dimensional torus
+$L^{p}$ real Lebesgue spaces
+$L^{p}_{\mathbb{C}}$ complex Lebesgue spaces
+$H^{1}$ real Sobolev space $W^{1,2}$
+$H^{1}_{\mathbb{C}}$ complex Sobolev space $W^{1,2}$, Eq. (6)
+$H^{-1}$ topological dual of the real Sobolev space
+$\mathcal{H}$ antisymmetric Hilbert space for $N$ particles, Eq. (5)
+$\mathscr{W}$ antisymmetric wave function space based on $H^{1}_{\mathbb{C}}$, Eq. (8)
+$\mathscr{D}$ set of density matrices, Eq. (19)
+$\mathscr{I}$ set of physical densities, Eq. (12)
+$\mathscr{X}$ affine space of normalized densities, Eq. (13)
+$\mathscr{X}_{>0}$ strictly positive subset of the affine space, Eq. (3a)
+$\mathscr{X}^{*}$ topological dual of the affine space, Eq. (3b)
+$\Omega^{\beta}_{v}$ Helmholtz free energy of a density matrices, Eq. (17)
+$\Omega^{\beta}$ Helmholtz functional on potentials, Eq. (18)
+$F^{\beta}_{\mathrm{DM}}$ thermal universal functional, Eq. (22b)
 II One-dimensional torus setting
 We consider a system at an elevated temperature $T$ consisting of $N$ spin $\frac{1}{2}$-particles.
+Choosing a different spin as an internal degree of freedom (or none at all) will not change any of the following results, we just use spin $\frac{1}{2}$ and antisymmetric wave functions to relate more closely to the usual systems considered in quantum chemistry.
 The particles are on the one-dimensional torus $\mathbb{T}$, meaning that they are confined to the interval $[0,1]$ where the points 0 and $1$ are identified with each other.
-The underlying Hilbert space is then
-$\mathcal{H}=\bigl{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}.$ (2)
-As noted, the considered Hamiltonians are of the form $H_{v}=H_{0}+V$ where $H_{0}$ consists of the kinetic energy operator and an interaction term and where $V$ is the potential.
-We can generally write $H_{v}$ instead of $H_{[v]}$, even though the considered potentials come from equivalence classes, since adding a constant to the potential just shifts the eigenvalues while not affecting the eigenstates.
-The interaction is kept general but positive and it needs to fulfill the KLMN condition (Reed and Simon [17], Th.
-X.17) with respect to the kinetic energy operator, i.e.
-for each interaction $W$ there are $0\leq a<1$ and $b\geq 0$ such that the following holds for each element in the form domain $Q(-\frac{1}{2}\Delta)$
+The underlying complex Hilbert space is then
+$\mathcal{H}=\bigl(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}.$ (5)
+We also define the complex Sobolev space on the torus
+$H_{\mathbb{C}}^{1}(\mathbb{T})=\{f\in L^{2}_{\mathbb{C}}(\mathbb{T})\mid\nabla f\in L^{2}_{\mathbb{C}}(\mathbb{T})\}$ (6)
+(like in Eq. (3a) this means a weak derivative) and the form domain of the kinetic energy operator
+$Q(-\tfrac{1}{2}\Delta)=\bigl(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}.$ (7)
+Together with the normalization constraint this gives the set of wave functions
+$\mathscr{W}=\{\Psi\in(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\mid\|\Psi\|_{\mathcal{H}}=1\}.$ (8)
+The one-particle density $\rho_{\Psi}$ for any $\Psi\in\mathscr{W}$ is obtained by integrating out all but one of the spatial coordinates from the modulus squared of a wave function summed over all spin components,
+$\rho_{\Psi}(x)=N\sum_{\sigma_{1},\ldots,\sigma_{N}}\int\lvert\Psi(x\sigma_{1},x_{2}\sigma_{2},\dotsc,x_{N}\sigma_{N})\rvert^{2}\mathrm{d}x_{2}\dotsb\mathrm{d}x_{N}.$ (9)
+As already stated in Theorem 1, the considered Hamiltonians are of the form $H_{v}=H_{0}+V$, where $H_{0}$ consists of the kinetic energy operator and an interaction term.
+We generally write $H_{v}$ instead of $H_{[v]}$, even though the considered potentials come from equivalence classes, since adding a constant to the potential just shifts the eigenvalues while not affecting the eigenstates.
+The interaction is kept general and just needs to fulfill the KLMN condition [34, Th.
+X.17] with respect to the kinetic energy operator.
+This means that for a fixed interaction $W$ there are $0\leq a<1$ and $b\geq 0$ such that for each $\Psi\in Q(-\frac{1}{2}\Delta)$ it holds
 $|\langle\Psi,W\Psi\rangle|\leq a\langle\Psi,-\tfrac{1}{2}\Delta\Psi\rangle+b\langle\Psi,\Psi\rangle.$ (10)
-This condition guarantees that the Hamiltonian $H_{0}=-\tfrac{1}{2}\Delta+W$ remains self-adjoint with the same form domain $Q(H_{0})=Q(-\frac{1}{2}\Delta)$ if we add an interaction to the kinetic energy operator (which is self-adjoint).
+This condition guarantees that the Hamiltonian $H_{0}=-\frac{1}{2}\Delta+W$ is self-adjoint with the same quadratic form domain as the kinetic energy operator.
 In Sutter et al.
-[7, Appendix] this condition was shown for interactions in $L^{1}([-1,1])$ and for distributional interactions of the form $\nabla g$ with $g\in L^{2}([-1,1])$.
-Here and in the following we use the notion of quadratic forms [18, Sec.
-VIII.6] instead of linear operators.
-For the kinetic energy operator the form domain is given by $Q(-\frac{1}{2}\Delta)=\bigl{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}$ with the Sobolev space $H^{1}(\mathbb{T})=\{\Psi\in L^{2}(\mathbb{T})\mid\nabla\Psi\in L^{2}(\mathbb{T})\}$ (like Eq.
-1a this means a weak derivative).
-We get by integration by parts
-$\frac{1}{2}\langle\Psi,-\Delta\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\Psi,-\Delta_{j}\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\nabla_{j}\Psi,\nabla_{j}\Psi\rangle.$ (4)
-Together with the normalization constraint this gives the set of wave functions
-$\mathscr{W}=\{\Psi\in(H^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\mid\|\Psi\|_{\mathcal{H}}=1\}.$ (5)
-The one-particle density $\rho_{\Psi}$ is obtained by integrating out all but one of the spatial coordinates of a wave function squared as well as summing over all spin components, i.e.
-$\rho_{\Psi}(x)=N\sum_{\sigma_{1},\ldots,\sigma_{N}}\int\lvert\Psi(x\sigma_{1},x_{2}\sigma_{2},\dotsc,x_{N}\sigma_{N})\rvert^{2}\mathrm{d}x_{2}\dotsb\mathrm{d}x_{N}.$ (6)
-In the following theorem we collect results from [6] and [7].
+[38, Appendix] this condition was shown to hold for interactions in $L^{1}([-1,1])$ and for distributional interactions of the form $\nabla g$ with $g\in L^{2}([-1,1])$.
+Here and in the following we use the notion of quadratic forms instead of linear operators [35, Sec.
+VIII.6].
+For the kinetic energy operator this is achieved directly by integration by parts
+$\frac{1}{2}\langle\Psi,-\Delta\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\Psi,-\Delta_{j}\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\nabla_{j}\Psi,\nabla_{j}\Psi\rangle=T(\Psi).$ (11)
+In the following theorem we collect some important previous results [28; 38].
 Theorem 2.
-We define the following set of densities
-$\mathscr{I}=\{\rho\in L^{1}(\mathbb{T})\mid\nabla\sqrt{\rho}\in L^{2}(\mathbb{T}),\rho\geq 0,\smallint\rho=N\}.$ (7)
+We define the following set of physical densities,
+$\mathscr{I}=\{\rho\in L^{1}(\mathbb{T})\mid\nabla\sqrt{\rho}\in L^{2}(\mathbb{T}),\rho\geq 0,\smallint\rho=N\},$ (12)
+and an affine space of normalized densities in the real Sobolev space $H^{1}(\mathbb{T})$,
+$\mathscr{X}=\{\rho\in H^{1}(\mathbb{T})\mid\smallint\rho=N\}.$ (13)
 Then, the following holds.
@@
   2. (ii)
-    There exist two constants $C_{1},C_{2}>0$ such that for all $\rho\in\mathscr{I}$ there is a $\Psi\in\mathscr{W}$ that has $\rho_{\Psi}=\rho$ and the following estimate holds
+    There exist two constants $C_{1},C_{2}>0$ such that for every $\rho\in\mathscr{I}$ there is a $\Psi\in\mathscr{W}$ that has $\rho_{\Psi}=\rho$ and fulfills the following estimate,
     $T(\Psi)\leq C_{1}+C_{2}\int\left|\nabla\sqrt{\rho(x)}\right|^{2}\mathrm{d}x.$ (15)
   3. (iii)
-    $\mathscr{X}_{>0}\subset\mathscr{I}\subset\mathscr{X}\subset H^{1}(\mathbb{T})\hookrightarrow C^{0}(\mathbb{T})\hookrightarrow L^{\infty}(\mathbb{T})\hookrightarrow L^{3}(\mathbb{T})\hookrightarrow L^{2}(\mathbb{T})\hookrightarrow L^{1}(\mathbb{T})$, where $\mathscr{X}=\{\rho\in H^{1}(\mathbb{T})\mid\smallint\rho=N\}$ and where $X\hookrightarrow Y$ means that $X$ is continuously embedded in $Y$.
+    $\mathscr{X}_{>0}\subset\mathscr{I}\subset\mathscr{X}\subset H^{1}(\mathbb{T})\hookrightarrow C^{0}(\mathbb{T})\hookrightarrow L^{\infty}(\mathbb{T})\hookrightarrow L^{3}(\mathbb{T})\hookrightarrow L^{2}(\mathbb{T})\hookrightarrow L^{1}(\mathbb{T})$, where $X\hookrightarrow Y$ means that $X$ is continuously embedded in $Y$.
 Lieb [28] showed that the set $\mathscr{I}$ is convex.
 This implies that (finite) convex combinations of elements in $\mathscr{I}$ lie again in $\mathscr{I}$.
-Motivated by the observations for the zero temperature case [7] we identify the space of external potential by $\mathscr{X}^{*}$, the topological dual of $\mathscr{X}$.
-This space contains equivalence classes, where the elements are from the space $H^{-1}(\mathbb{T})$ and where two elements are in the same class if they differ only by a constant, i.e., like in Eq.
+Motivated by the observations for the zero temperature case [38], we identify the space of external potential by $\mathscr{X}^{*}$, the topological dual of the subspace $\{\rho\in H^{1}(\mathbb{T})\mid\int\rho=0\}$ that is isomorphic to $\mathscr{X}$.
+This space $\mathscr{X}^{*}$ contains equivalence classes, where the elements are from the space $H^{-1}(\mathbb{T})$ and where two elements are in the same class if and only if they differ only by a constant, i.e., like in Eq.
 (3b),
@@
 The dual pairing of an element $v=f+\nabla g\in H^{-1}(\mathbb{T})$ with an element $\varphi\in H^{1}(\mathbb{T})$ is given by $v(\varphi)=\langle v,\varphi\rangle=\langle f,\varphi\rangle-\langle g,\nabla\varphi\rangle$.
-Using the dual space of $\mathscr{X}$ for the external potentials guarantees that the potential energy is finite.
-That Hamiltonians of the form $H=H_{0}+V$ with $V$ defined by a $v\in\mathscr{X}^{*}$ are connected to a self-adjoint operator is shown in [7] by using the KLNM theorem (Reed and Simon [17], Th.
-X.17).
+Using $\mathscr{X}^{*}$ as potential space automatically guarantees that the potential energy is finite.
+That Hamiltonians of the form $H=H_{0}+V$ with $V$ defined by a $v\in\mathscr{X}^{*}$ are connected to a self-adjoint operator was shown in Sutter et al.
+[38, Th.
+18] by using the KLMN theorem [34, Th.
+X.17].
 III Helmholtz free energy and Gibbs state
 Considering a system at fixed temperature $T>0$ (through coupling to a heat bath) means that we cannot rely on a variational principle for the internal energy $E_{v}(\Gamma)=\operatorname{Tr}\{\Gamma H_{v}\}$ any more [3].
-Instead, we have to consider the Helmholtz free energy
+Instead, we have to consider the Helmholtz free energy for arbitrary density matrices,
 $\Omega_{v}^{\beta}(\Gamma)=E_{v}(\Gamma)-TS(\Gamma)=\operatorname{Tr}\bigl\{\Gamma\bigl(H_{v}+\beta^{-1}\log(\Gamma)\bigr)\bigr\},$ (17)
-where $S(\Gamma)=-\operatorname{Tr}\{\Gamma\log(\Gamma)\}$ is the entropy and $\beta=T^{-1}$ is the inverse temperature. The new variational principle then yields the Helmholtz functional
-$\Omega^{\beta}(v)=\inf_{\Gamma}\Omega^{\beta}_{v}(\Gamma)=\inf_{\Gamma}\operatorname{Tr}\bigl{\{}\Gamma\bigl{(}H_{v}+\beta^{-1}\log(\Gamma)\bigr{)}\bigr{\}}$ (12)
-and can been equivalentely seen as the Legendre transform of $E_{v}$ (in terms of entropy) to $\Omega^{\beta}(v)$ (in terms of temperature).
-Here the infimum is over all density matrices, i.e.
-over all (possibly infinite) convex sums $\Gamma=\sum_{k}\lambda_{k}\langle\Psi_{k},\cdot\rangle\Psi_{k}$, $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$, with $\Psi_{k}\in\mathscr{W}$ all orthogonal.
-Making the Helmholtz functional stationary with respect to small perturbation $\delta\Gamma$ with $\operatorname{Tr}\{\delta\Gamma\}=0$ gives the minimizer.
+where $S(\Gamma)=-\operatorname{Tr}\{\Gamma\log(\Gamma)\}$ is the entropy and $\beta=T^{-1}$ is the inverse temperature. The variational principle then yields the Helmholtz functional,
+$\Omega^{\beta}(v)=\inf_{\Gamma\in\mathscr{D}}\Omega^{\beta}_{v}(\Gamma)=\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\bigl\{\Gamma\bigl(H_{v}+\beta^{-1}\log(\Gamma)\bigr)\bigr\}.$ (18)
+Here the infimum is taken over the set of all density matrices,
+$\mathscr{D}=\left\{\Gamma=\sum_{k=1}^{\infty}\lambda_{k}\langle\Psi_{k},\cdot\rangle\Psi_{k}\;\middle|\;\lambda_{k}\geq 0,\operatorname{Tr}\{\Gamma\}=1,\Psi_{k}\in\mathscr{W}\;\text{all orthogonal},S(\Gamma)<\infty\right\}.$ (19)
+Making the Helmholtz free energy stationary with respect to small perturbations $\delta\Gamma$ with $\operatorname{Tr}\{\delta\Gamma\}=0$ gives the minimizer.
 Proposition 3.
-The unique minimizer of the Helmholtz functional is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ where $Z(v)=\operatorname{Tr}\{\mathrm{e}^{-\beta H_{v}}\}$ is called the partition function.
-The state $\Gamma_{v}$ is called the Gibbs state and it does only exist if $Z(v)$ is finite.
+The unique minimizer of the Helmholtz free energy is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)\in\mathscr{D}$, where $Z(v)=\operatorname{Tr}\{\mathrm{e}^{-\beta H_{v}}\}$ is called the partition function.
+The state $\Gamma_{v}$ is called the Gibbs state and only exists if $Z(v)$ is finite.
 Moreover, for each $v\in H^{-1}(\mathbb{T})$ the Gibbs state exists and the Helmholtz free energy is given by
-$\Omega^{\beta}(v)=-\beta^{-1}\log\bigl{(}Z(v)\bigr{)}.$ (13)
-The proof of Proposition 3 can be found in Appendix A.
-The reason for $Z(v)$ to be finite is that the particles are trapped in a finite volume.
-The defining parameters of canonical (quantum) thermodynamics are $(N,V,T)$ and all of these parameters need to be finite.
-For a counterexample [16, Sec.
-2.3] consider the hydrogen atom in $\mathbb{R}^{3}$.
-The energies are $1/(2n^{2})$ with a degeneracy of $n^{2}$ and with $n\in\mathbb{N}$.
+$\Omega_{v}^{\beta}(\Gamma_{v})=\Omega^{\beta}(v)=-\beta^{-1}\log\bigl(Z(v)\bigr).$ (20)
+Remark 1.
+For an example of a physical system with infinite partition function simply consider the hydrogen atom in the infinite-volume space $\mathbb{R}^{3}$ [17, Sec.
+2.3].
+The eigenenergies are $1/(2n^{2})$ with multiplicity $^{2}, nn\in\mathbb{N}$.
 Then the partition function is
 $Z(v)=\sum\limits_{n=1}^{\infty}n^{2}\mathrm{e}^{\beta/(2n^{2})}\geq\sum\limits_{n=1}^{\infty}n^{2}=\infty.$ (21)
+The proof of Proposition 3 can be found in Appendix A.
 Proposition 3 implies that the density of the Gibbs state is an infinite convex combination of elements in $\mathscr{I}$.
@@
 Proposition 4.
-The density of the Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$ for all $v\in H^{-1}$.
+The density of the Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$ for all $v\in H^{-1}(\mathbb{T})$.
 The proposition is proven in Appendix B.
-It tells us that $v$-representable densities at elevated temperatures are contained in $\mathscr{I}$.
-Hence we can split the minimization of the Helmholtz functional $\Omega^{\beta}_{v}$ as in the zero temperature case into first a variation over densities in $\mathscr{I}$ followed by a minimization over all density matrices which yield such a density,
+It tells us that $v$-representable densities at elevated temperatures are always contained in $\mathscr{I}$.
+Hence we can split the minimization of the Helmholtz free energy $\Omega^{\beta}_{v}$ into first a variation over densities in $\mathscr{I}$ followed by a minimization over all density matrices which yield such a density, just like in the zero temperature case,
 $\Omega^{\beta}(v)=\inf_{\rho\in\mathscr{I}}\{F^{\beta}_{\mathrm{DM}}(\rho)+\langle v,\rho\rangle\}\quad\text{where}$ (22a)
-$F^{\beta}_{\mathrm{DM}}(\rho)=\inf_{\Gamma\mapsto\rho}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}.$ (15b)
-The functional $F_{\mathrm{DM}}^{\beta}(\rho)$ is called the universal functional and we can extend it to $\mathscr{X}$ by setting $F_{\mathrm{DM}}^{\beta}(\rho)=\infty$ for all $\rho\in\mathscr{X}\setminus\mathscr{I}$.
-Note that because of the existence of the Gibbs state, the infimum in the variation over the densities is attained whereas for the second infimum (in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$) we only know about the existence of a minimizer for $v$-representable densities which are characterized in Theorem 1.
-As for zero temperature ([6; 7]), we would like to have a convex formulation of thermal density-functional theory to exploit the properties of convex functions.
-Due to the concavity of the entropy [16, Th.
+$F^{\beta}_{\mathrm{DM}}(\rho)=\inf_{\begin{subarray}{c}\Gamma\in\mathscr{D}\\ \Gamma\mapsto\rho\end{subarray}}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}.$ (22b)
+The functional $F_{\mathrm{DM}}^{\beta}(\rho)$ is called the thermal universal functional and we can extend it to $\mathscr{X}$ by setting $F_{\mathrm{DM}}^{\beta}(\rho)=\infty$ for all $\rho\in\mathscr{X}\setminus\mathscr{I}$.
+Note that because of the existence of the Gibbs state and Propositions 3 and 4, the infimum in the variation over the densities is always attained, whereas for the second infimum (in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$) we only know about the existence of a minimizer for $v$-representable densities which are characterized in Theorem 1.
+Like in the zero-temperature case [28; 38], we would like to have a convex formulation of thermal density-functional theory to exploit the properties of convex functions.
+Due to the concavity of the entropy [17, Th.
 12] and linearity of $\Gamma\to\rho$ the following proposition follows.
@@
   1. (i)
-    The Helmholtz functional $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)$ is convex.
+    The Helmholtz free energy $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)$ is convex on $\mathscr{D}$.
   2. (ii)
-    The universal functional $\rho\mapsto F^{\beta}_{\mathrm{DM}}(\rho)$ is convex on $\mathscr{X}$.
-IV Relation between $v$-representability and subdifferentials
-The relation between $v$-representability and non-empty subdifferential was presented in the literature [20; 21; 22].
-This relation, together with some additional results from convex analysis, are important ingredients to show our main result Theorem 1.
+    The thermal universal functional $\rho\mapsto F^{\beta}_{\mathrm{DM}}(\rho)$ is convex on $\mathscr{X}$.
+IV Relation between $v$-representability and non-empty subdifferential
+The relation between $v$-representability and non-empty subdifferential is usually discussed in mathematical presentations of DFT [12; 32; 6].
+This relation, together with some additional results from convex analysis, are important ingredients to show our main Theorem 1.
 Therefore, we will repeat the relation and the needed results in this section.
@@
 An element of $\partial f(\rho)$ is called a subgradient.
-Moreover, the function $f$ is Gâteaux differentiable at a point $\rho$ if the function $f$ is continuous at $\rho$ and if the subdifferential $\partial f(\rho)$ contains only a single element [23, Prop.
+Moreover, the function $f$ is Gâteaux differentiable at a point $\rho$ if the function $f$ is continuous at $\rho$ and if the subdifferential $\partial f(\rho)$ contains only a single element [2, Prop.
 2.40].
 In this case, this element is equal to the derivative.
-If the general convex function is replaced by the universal functional (which is a proper convex functional by Proposition 5), we get for an element $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ and for all $\rho^{\prime}\in\mathscr{X}$ the following inequality
+If the general convex function is replaced by the thermal universal functional (which is a proper convex functional by Proposition 5), we get for an element $v\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)$ and for all $\rho^{\prime}\in\mathscr{X}$ the following inequality
 $F_{\mathrm{DM}}^{\beta}(\rho)+\langle v,\rho\rangle\leq F_{\mathrm{DM}}^{\beta}(\rho^{\prime})+\langle v,\rho^{\prime}\rangle.$ (24)
@@
 In the next section we then show that these densities indeed come from Gibbs states.
-The two result together give then $v$-representability for all densities in $\mathscr{X}_{>0}$.
+Only the two result together then give $v$-representability for all densities in $\mathscr{X}_{>0}$.
 To show the existence of a subgradient for a density $\rho\in\mathscr{X}_{>0}$ we proceed as in Sutter et al.
 [38].
-The only result that significantly changes compared to this reference is the local boundedness of the universal functional $F_{\mathrm{DM}}^{\beta}$.
+The only result that significantly changes compared to our previous work is the local boundedness of the thermal universal functional $F_{\mathrm{DM}}^{\beta}$.
 Lemma 6.
@@
 For the lower bound of $F_{\mathrm{DM}}^{\beta}(\rho)$ we have
-$F_{\mathrm{DM}}^{\beta}(\rho)\geq\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}=\Omega^{\beta}(0)$ (19)
-which is greater than -$\infty$ by Proposition 3. For the upper bound we note that the entropic term $\beta S(\Gamma)$ is negative. Hence we get the following inequality
-$F_{\mathrm{DM}}^{\beta}(\rho)\leq\inf_{\Gamma\to\rho}\operatorname{Tr}\{\Gamma H_{0}\}\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2}$ (20)
-where we used a previous result [7, Lem. 10] for the second inequality. ∎
+$F_{\mathrm{DM}}^{\beta}(\rho)\geq\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}=\Omega^{\beta}(0),$ (26)
+which is greater than -$\infty$ by Proposition 3. For the upper bound we note that the entropic term $\beta^{-1}S(\Gamma)$ is negative. Hence we get the following inequality
+$F_{\mathrm{DM}}^{\beta}(\rho)\leq\inf_{\begin{subarray}{c}\Gamma\in\mathscr{D}\\ \Gamma\mapsto\rho\end{subarray}}\operatorname{Tr}\{\Gamma H_{0}\}\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2},$ (27)
+where we used a previous result [38, Lem. 10] for the second inequality. ∎
 Theorem 7.
-For any $\rho\in\mathscr{X}_{>0}$ the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous and the subdifferential $\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$ is non-empty.
-The proof is the same as in Sutter et al. [7, Th. 15].
+At any $\rho\in\mathscr{X}_{>0}$ the thermal universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous and the subdifferential $\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$ is non-empty.
+The proof is the same as in Sutter et al.
+[38, Th.
+15].
+The result relies critically on the embedding $H^{1}\hookrightarrow L^{\infty}$ from Theorem 2(iii) that can be used to show that $\mathscr{X}_{>0}$ is an open set [38, Lem.
+14].
+This embedding only holds true in one dimension and one further needs the torus setting with a non-vanishing minimum of $\rho$ in order to show that $F_{\mathrm{DM}}^{\beta}$ is bounded above.
+Both results combined then yield the desired result by using powerful theorems from convex analysis [2, Th.
+2.14 and Prop.
+2.36].
+We note that the same result can be established also on the open interval with Neumann boundary conditions [8].
+We can now also understand the origin of the distributional potentials.
+We have $\mathscr{X}_{>0}\subset\mathscr{X}\subset H^{1}(\mathbb{T})$, so the density domain is limited to a subset of the Sobolev space $H^{1}(\mathbb{T})$ with dual space $H^{-1}(\mathbb{T})\subset\mathscr{X}^{*}$ that contains distributions.
+On the other hand, a potential is always from the subdifferential of the thermal universal functional and the subdifferential itself is a subset of the dual space $\mathscr{X}^{*}$.
 V Existence of the Minimum
@@
 (22b)) exists.
-This minimizer must then be the Gibbs state $\Gamma_{v}$, where $v$ is just the potential that yields $\rho$.
+This minimizer is then the Gibbs state $\Gamma_{v}$, where $v$ is just the potential that yields $\rho$.
 In the following we show this for all $\rho\in\mathscr{X}_{>0}$.
-To do so, we make use of the relative entropy which is defined as
-$S(A|B)=\operatorname{Tr}\{A(\log A-\log B)+B-A\}$ (21)
-for two non-negative trace class operators $A$, $B$ with $\ker B\subset\ker A$. An important property of the relative entropy is lower semi-continuity.
+To do so, we make use of the relative entropy which for two non-negative trace class operators $A$, $B$ with $\ker B\subset\ker A$ is defined as
+$S(A|B)=\operatorname{Tr}\{A(\log A-\log B)+B-A\}.$ (28)
+The relative entropy is always non-negative, $S(A|B)\geq 0$ and $S(A|B)=0$ if and only if $A=B$ [29; 14]. An important property of the relative entropy is its lower semi-continuity.
 Lemma 8 (Lami and Shirokov [24], Lemma 4).
-The relative entropy $(A,B)\mapsto S(A|B)$ is jointly lower semi-continuous with respect to the weak-* topology. (for non-negative trace class operators).
-The next lemma relates the Helmholtz functional and the relative entropy.
-Note that originally we only defined the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)$ for (normalized) density matrices.
-However, in the same manner it can be defined for any non-negative trace class operator.
-Recall that $\Gamma_{v}$ is the Gibbs state corresponding to a potential $v$ and it is given by $\mathrm{e}^{-\beta H_{v}}/Z(v)$.
+The relative entropy $(A,B)\mapsto S(A|B)$ is jointly lower semi-continuous with respect to the weak-* topology.
+The next lemma relates the Helmholtz free energy and the relative entropy.
+Note that originally we only defined the Helmholtz free energy $\Omega^{\beta}_{v}(\Gamma)$ for $\Gamma\in\mathscr{D}$.
+However, in the same manner, it can be defined for any non-negative trace-class operator with finite entropy.
+Recall that $\Gamma_{v}$ is the Gibbs state corresponding to a potential $v$ and is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$.
+Its kernel is trivial and thus $S(\cdot|\Gamma_{v})$ is well-defined.
 Lemma 9 (Giesbertz and Ruggenthaler [17]).
 $S(\Gamma|\Gamma_{v})+\operatorname{Tr}\{\Gamma-\Gamma_{v}\}=\beta\Omega^{\beta}_{v}(\Gamma)-\beta\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}$.
-Since the left hand side in Lemma 9 consists of two weak-* lower semi-continuous functions (norm is decreasing under weak-* limits), we get that the Helmholtz functional together with a term proportional to $\operatorname{Tr}\{\cdot\}$ is lower semi-continuous with respect to the weak-* topology.
+Note that the left-hand side is well-defined for all non-negative trace-class operators $\Gamma$ and thus Lemma 9 extends the domain of $\Omega_{v}^{\beta}(\cdot)$ to density matrices with infinite entropy.
+Since the left-hand side in Lemma 9 consists of two weak-* lower semi-continuous functions (the trace norm $\operatorname{Tr}|\Gamma|$ is decreasing under weak-* limits and equal to $\operatorname{Tr}\{\Gamma\}$ since $\Gamma\geq 0$), we get that the Helmholtz free energy together with a term proportional to $\operatorname{Tr}\{\cdot\}$ (with positive factor) is lower semi-continuous with respect to the weak-* topology.
 In particular, we get the following lemma.
@@
 Theorem 11.
-For any $\rho\in\mathscr{X}_{>0}$ there is a (normalized) density matrix $\Gamma\mapsto\rho$ such that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log\Gamma)\}$.
-This shows the existence of a minimizer in Eq.
-(15b).
+For every $\rho\in\mathscr{X}_{>0}$ and all potentials $v\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$, it holds that the corresponding Gibbs state $\Gamma_{v}$ yields the density $\rho$ and is the minimizer in Eq.
+(22b), i.e., $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{v}(H_{0}+\beta^{-1}\log\Gamma_{v})\}$.
 Proof.
-For a density $\rho\in\mathscr{X}_{>0}$ we pick $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ which we know to exists by Theorem 7.
+For a density $\rho\in\mathscr{X}_{>0}$ we pick $v\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)$ which we know to exist by Theorem 7.
 By definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ there is a sequence of density matrices $\Gamma_{n}$ each yielding the density $\rho$ and such that $\operatorname{Tr}\{\Gamma_{n}(H_{0}+\beta^{-1}\log\Gamma_{n})\}\to F^{\beta}_{\mathrm{DM}}(\rho)$.
-Since all $\Gamma_{n}$ are contained in a ball of radius $1$ we get by the Banach–Alaoglu theorem [18, Th.
+Since all $\Gamma_{n}$ are contained in a ball of radius $1$ we get by the Banach–Alaoglu theorem [35, Th.
 IV.21] that there is a subsequence (in the following also denoted by $\Gamma_{n}$) and a trace class operator $\Gamma_{0}$ such that $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$.
 Moreover, we know that $\Gamma_{0}$ is non-negative and that $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$.
-To prove the theorem, we show that $\Gamma_{0}$ is a minimizer.
-That means that $\Gamma_{0}$ is normalized to $1$, that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ and that it yields the density $\rho$.
+To prove the theorem, we show that $\Gamma_{0}$ is a minimizer in Eq.
+(22b) that, together with the discussion after Eq.
+(24), already shows that $\Gamma_{0}$ is equal to the Gibbs state $\Gamma_{v}\in\mathscr{D}$.
+This means that $\Gamma_{0}$ is normalized to $1$, that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$, and that it yields the density $\rho$.
 We start by showing $\operatorname{Tr}\{\Gamma_{0}\}=1$.
-For simplicity, we call the minimum of the Helmholtz functional $\Omega^{\beta}=\Omega^{\beta}_{v}(\Gamma_{v})=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$.
-Because $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$ and $\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}\}=1$ we have
+For simplicity, we call the minimum of the Helmholtz free energy $\Omega^{\beta}=\Omega^{\beta}_{v}(\Gamma_{v})=\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$.
+Because $v\in-\partial F^{\beta}_{\mathrm{DM}}(\rho)$ and $\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}\}=1$ we have
 $\Omega^{\beta}-\xi\Omega^{\beta}$ $=F^{\beta}_{\mathrm{DM}}(\rho)+\langle v,\rho\rangle-\xi\Omega^{\beta}=\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}(H_{v}+\beta^{-1}\log\Gamma_{n})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{n}\}$ (30)
-$\geq\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\}$
+$\geq\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\},$
 where we have used the fact that the function $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous due to Lemma 10.
@@
 This expression is well-defined, since we can choose a representative of $v]\in\mathscr{X}^{*}$ for which $\Omega^{\beta}_{v}(\Gamma_{v})$ is non-zero.
-So the condition of Lemma 10 is satisfied.
 Let $c$ be the constant such that $\operatorname{Tr}\{c\Gamma_{0}\}=1$.
 (Note that the constant $c$ only exist if $\Gamma_{0}\neq 0$.
-If $\Gamma_{0}=0$ then the right hand-side of Eq.
-(23) is 0 where the left hand-side is $\beta^{-1}$.
+If $\Gamma_{0}=0$ then the right-hand side of Eq.
+(30) is 0 while the left-hand side is $\beta^{-1}$.
 Hence $\Gamma_{0}=0$ is not possible.) Since $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$ it holds $c\geq 1$.
-Then we can compute
+We compute
 $\Omega^{\beta}-\xi\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})-\xi\Omega^{\beta}\operatorname{Tr}\{c\Gamma_{0}\}$ $=\operatorname{Tr}\{c\Gamma_{0}(H_{v}+\beta^{-1}\log(c\Gamma_{0}))\}-\xi\Omega^{\beta}\operatorname{Tr}\{c\Gamma_{0}\}$ (31)
@@
 $=c(\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\})+\beta^{-1}\log c$
-$\leq c(\Omega^{\beta}-\xi\Omega^{\beta})+\beta^{-1}\log c$
-where we used the variational principle for $\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})$, the fact that $\operatorname{Tr}\{c\Gamma_{0}\}=1$, and Eq. (23). This inequality can be transformed to
+$\leq c(\Omega^{\beta}-\xi\Omega^{\beta})+\beta^{-1}\log c,$
+where we used the variational principle for $\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})$, the fact that $\operatorname{Tr}\{c\Gamma_{0}\}=1$, and Eq.
+(30).
+This inequality can be transformed into
 $c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}\geq 0.$ (32)
-It is clear that we have equality if $c$ is equal to $1$.
+It is clear that we have equality if $c=1$.
 The function $f(c)=c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}$ defined on $1,\infty)$ has $f(1)=0$ and the derivative
@@
 Hence the inequality (32) can only be satisfied if $c=1$ and we conclude that $\operatorname{Tr}\{\Gamma_{0}\}=1$.
-Now that we know that the weak-* limit $\Gamma_{0}$ has trace norm equal to $1$ we can show that $\Gamma_{0}\to\rho$.
-For this we use the same idea as Lieb [6] and show that for any $f\in L^{\infty}(\mathbb{T})$ we have $\operatorname{Tr}\{\Gamma_{n}M_{f}\}\to\operatorname{Tr}\{\Gamma_{0}M_{f}\}$ where $M_{f}$ denotes pointwise multiplication by $f$.
-This means that the densities of $\Gamma_{n}$ converge weakly to the density of $\Gamma_{0}$ (in $L^{1}$) and by the uniqueness of weak limits we get $\Gamma_{0}\to\rho$.
+Now that we know that the weak-* limit $\Gamma_{0}$ has trace norm equal to $1$ we can show that $\Gamma_{0}\mapsto\rho$.
+For this we use the same idea as Lieb [28] and show that for any $f\in L^{\infty}(\mathbb{T})$ we have $\operatorname{Tr}\{\Gamma_{n}M_{f}\}\to\operatorname{Tr}\{\Gamma_{0}M_{f}\}$, where $M_{f}$ denotes pointwise multiplication by $f$.
+This means that the densities of $\Gamma_{n}$ converge weakly to the density of $\Gamma_{0}$ (in $L^{1}$) and by the uniqueness of weak limits we get $\Gamma_{0}\mapsto\rho$.
 To do this, let $|\sum_{j}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of $\Gamma_{0}$.
@@
 $|.P=\sum\limits_{j=1}^{N_{\epsilon}}\ket{\psi_{j}}\bra{\psi_{j}}$. (35)
-The operator $P$ is compact and since $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$ in trace class, there exist $N_{P}$ such that $\lvert\operatorname{Tr}\{\Gamma_{n}P\}-\operatorname{Tr}\{\Gamma_{0}P\}\rvert<\epsilon$ for all $n>N_{P}$.
+The operator $P$ is compact and since $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$ is trace class there exists an $N_{P}$ such that $\lvert\operatorname{Tr}\{\Gamma_{n}P\}-\operatorname{Tr}\{\Gamma_{0}P\}\rvert<\epsilon$ for all $n>N_{P}$.
 It follows that for $n>N_{P}$ we have $\operatorname{Tr}\{\Gamma_{n}P\}>\operatorname{Tr}\{\Gamma_{0}P\}-\epsilon>1-2\epsilon$ and thus $\operatorname{Tr}\{\Gamma_{n}(1-P)\}<2\epsilon$.
-Then we get that $\lvert\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)\}\rvert<2\epsilon$ for large enough $n$.
-We can write
+We thus get that $\lvert\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)\}\rvert<2\epsilon$ for large enough $n$ and we can write
 $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}=\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})PM_{f}\}+\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)M_{f}\}.$ (36)
@@
 The second term is contained in $C\epsilon,C\epsilon]$ for $n$ large enough where the constant $C$ depends only on $f$.
-Since $\epsilon$ can be arbitrarily small we conclude that $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}\to 0$ and hence $\Gamma_{0}\to\rho$.
+Since $\epsilon$ can be arbitrarily small we conclude that $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}\to 0$ and hence $\Gamma_{0}\mapsto\rho$.
 By Eq.
@@
 Note that we never use $\rho>0$ and that we can also replace the condition $\rho\in\mathscr{X}_{>0}$ by the condition that the subdifferential of $F_{\mathrm{DM}}^{\beta}$ at $\rho$ is non-empty.
-If we combine Theorem 7 and Theorem 11, we get that each density $\rho\in\mathscr{X}_{>0}$ is $v$-representable, which is the first part of Theorem 1.
+If we combine Theorem 7 and Theorem 11, we get that $\rho\in\mathscr{X}_{>0}$ is a sufficient condition for $v$-representability by a Gibbs state, which is the first part of Theorem 1.
 The full proof for Theorem 1 can be found in Section VIII.
 VI Closing the Circle
-So far we showed that each density in $\mathscr{X}_{>0}$ is $v$-representable.
-In this section we show that each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$.
-At elevated temperatures, all excited state of the Hamiltonian contribute to the Gibbs state (and hence also to the density) and since these state form an orthonormal basis, it is not possible that the density vanishes somewhere.
+So far we showed that each density in $\mathscr{X}_{>0}$ is $v$-representable by a Gibbs state.
+In this section we show that each $v$-representable density is strictly positive and thus contained in $\mathscr{X}_{>0}$.
+Consequently, the restriction to strictly positive densities in order to show $v$-representability of a density is necessary.
+In order to show this, observe that at elevated temperatures, all excited states of the Hamiltonian contribute to the Gibbs state (and hence also to the density) and since these states form an orthonormal basis, it is not possible that the density vanishes somewhere.
 We formulate this result in the following theorem that closes the circle from strictly positive densities in $\mathscr{X}_{>0}$ to representing potentials in $\mathscr{X}^{*}$ and back to a strictly positive density of the corresponding Gibbs state.
 Theorem 12.
-The density of any Gibbs state is everywhere non-zero.
+The density of any Gibbs state is strictly positive.
 Proof.
@@
 $|,\Gamma_{v}=\frac{1}{Z(v)}\sum_{j}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$, (37)
-where the $\lambda_{j}$ are the eigenvalues of $H_{v}$ with corresponding eigenstates $\psi_{j}$.
+where the $\lambda_{j}$ are the eigenvalues of $H_{v}$ with corresponding eigenstates $\psi_{j}\in\mathscr{W}$.
 We denote the density of $\Gamma_{v}$ by $\rho$ and the densities of $\psi_{j}$ by $\rho_{j}$.
@@
 we conclude that $\rho_{j}(x_{0})=0$ for all $j$ since the weights are all non-zero.
-This implies that all $\psi_{j}$ vanish on $U=\{x_{0}\}\times[0,1]^{N-1}$.
-(The restriction of $\psi_{j}$ to $U$ is well-defined by the trace operator for Sobolev spaces [25, p.
-315].) Consequently, each linear combination of the $\psi_{j}$ vanishes on $U$.
-Yet, this statement makes only sense if the linear combination is in $H^{1}$ too.
-Since the $\psi_{j}$ form a basis of $L^{2}(\mathbb{T}^{N})$ we also need to be able to express each function in $H^{1}(\mathbb{T}^{N})\subset L^{2}(\mathbb{T}^{N})$ as their linear combination.
-Now this yields a contradiction since clearly not every function in $H^{1}(\mathbb{T}^{N})$ vanishes on U.
-We note that while the restriction to $U$ is well-defined for $H^{1}(\mathbb{T}^{N})$ functions, this is not the case for functions in $L^{2}(\mathbb{T}^{N})$.
+This implies that all $\psi_{j}$ vanish on $U=\{x_{0}\}\times[0,1]^{N-1}\times\{\uparrow,\downarrow\}^{N}$.
+Note that the restriction to $U$ of $\psi_{j}\in\mathscr{W}$ or any linear combination of those vectors in $(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ is well-defined by the trace operator for Sobolev spaces [4, p.
+315], which is not the case for a general element in $\mathcal{H}=(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$.
+Since the $\psi_{j}$ form a basis of $(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ we also need to be able to express each function in $(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\subset(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ as their linear combination.
+Now this yields a contradiction since clearly not every function in $(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}$ vanishes on $U$.
 ∎
@@
 The last part of the main result (Theorem 1) is about Gâteaux differentiability of $F_{\mathrm{DM}}^{\beta}$ at densities $\rho\in\mathscr{X}_{>0}$.
-As mentioned in Section IV, Gâteaux differentiability can be achieved by continuity and a single-element subdifferential.
+As mentioned in Section IV, Gâteaux differentiability can be achieved by continuity and a single-valued subdifferential.
 Thus, the goal of this section is to show that each density $\rho\in\mathscr{X}_{>0}$ has only one subgradient.
-We do that by using standard results from convex analysis and by exploiting the following lemma.
-Lemma 13.
-The Helmholtz free energy $\Omega^{\beta}(v)$ is a strictly concave function in $v\in H^{-1}(\mathbb{T})$.
-This result is obtained by exploiting the uniqueness of the Gibbs state. The proof proceeds analogously to [16, Th. 10]. Additionally, we need the following definition and lemmata.
-Definition 14 (Barbu and Precupanu [23], p. 71).
-Let $f$ be a convex function on $\mathscr{X}$.
-The closure of $f$, denoted by $\mathrm{cl}f$, is the lower semi-continuous hull of $f$, i.e.
-$\mathrm{cl}f(\rho)=\liminf_{\rho^{\prime}\to\rho}f(\rho^{\prime})$ for all $\rho\in\mathscr{X}$ if $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})>$-$\infty$ for all $\tilde{\rho}$ and it is the constant function -$\infty$ if there is a $\tilde{\rho}\in\mathscr{X}$ such that $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})=$-$\infty$.
-Since the universal functional $F_{\mathrm{DM}}^{\beta}$ is uniformly bounded from below by Lemma 6, the closure of it is given by $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$.
-Another characterization of the closure of a convex function is by its biconjugate as the following lemma shows.
-Lemma 15 (Barbu and Precupanu [23], Cor. 2.23).
-The biconjugate of a convex function $f$ coincides with its closure. That is, $f^{**}=\mathrm{cl}f$.
-Lemma 16 (Barbu and Precupanu [23], Prop. 2.33).
-Let $f:\mathscr{X}\to\mathbb{R}\cup\{\infty\}$ be a proper, convex and lower semi-continuous function and $f^{*}$ its conjugate. Then the following two statements are equivalent,
-  1. (i)
-    $v\in\partial f(\rho)$
-  2. (ii)
-    $\rho\in\partial f^{*}(v)$.
+We do that by using a standard result from convex analysis and by the usual Hohenberg–Kohn argument.
 Lemma 13 (Barbu and Precupanu [2], Prop. 2.40).
@@
 Conversely, if $f$ is continuous at $\rho$ and if $\partial f(\rho)$ contains only a single element, then $f$ is Gâteaux differentiable at $\rho$ and $\mathrm{grad}f(\rho)=\partial f(\rho)$.
-We have now all the needed ingredients to show Gâteaux differentiability of the universal functional.
+We now have all the necessary ingredients to show Gâteaux differentiability of the thermal universal functional.
 Theorem 14.
-The universal functional $F_{\mathrm{DM}}^{\beta}$ is Gâteaux differentiable at any density $\rho\in\mathscr{X}_{>0}$.
+The thermal universal functional $F_{\mathrm{DM}}^{\beta}$ is Gâteaux differentiable at any density $\rho\in\mathscr{X}_{>0}$.
 Proof.
-According to Theorem 7, the functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$.
-By Lemma 17 it is then enough to show that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued for all densities $\rho\in\mathscr{X}_{>0}$.
-The conjugate function of $F_{\mathrm{DM}}^{\beta}$ is given by $\Omega^{\beta}(-\cdot)$.
-If we conjugate this function we end up with the biconjugate of $F_{\mathrm{DM}}^{\beta}$ which is defined by
-$F_{\mathrm{DM}}^{\beta**}(\rho)=\sup\big{\{}\Omega^{\beta}(v)-\langle\rho,v\rangle\mid v\in H^{-1}(\mathbb{T})\big{\}}.$ (32)
-The biconjugate $F_{\mathrm{DM}}^{\beta**}$ is a convex, lower semi-continuous function and it coincides with the convex envelope of $F_{\mathrm{DM}}^{\beta}$.
-Since $F_{\mathrm{DM}}^{\beta}$ is convex, Lemma 15 states that the closure of $F_{\mathrm{DM}}^{\beta}$ equals to its biconjugate $F_{\mathrm{DM}}^{\beta**}$.
-By Definition 14 and its proceeding discussion $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$.
-Theorem 7 shows that the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$ and hence we get that for each $\rho>0$
-$F_{\mathrm{DM}}^{\beta**}(\rho)=\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})=F_{\mathrm{DM}}^{\beta}(\rho).$ (33)
-This means that $F_{\mathrm{DM}}^{\beta}$ and its biconjugate $F_{\mathrm{DM}}^{\beta**}$ coincide on the open set $\mathscr{X}_{>0}$.
-In particular this means that the subdifferentials are equal for densities $\rho\in\mathscr{X}_{>0}$, i.e.
-$\partial F_{\mathrm{DM}}^{\beta}(\rho)=\partial F_{\mathrm{DM}}^{\beta**}(\rho)$.
-So instead of showing that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued, we show it for $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$.
-Note that the conjugate function of $F_{\mathrm{DM}}^{\beta**}$ is $\Omega^{\beta}(-\cdot)$ since the triple conjugate of $F_{\mathrm{DM}}^{\beta}$ is equal to the conjugate.
-The function $F_{\mathrm{DM}}^{\beta**}$ fulfills all assumptions of Lemma 16 and hence we get
-$v\in\partial F_{\mathrm{DM}}^{\beta**}(\rho)\Longleftrightarrow\rho\in\partial\left(-\Omega^{\beta}\right)(v).$ (34)
-By Lemma 13 we have that $\Omega^{\beta}$ is strictly convex.
-In particular this means that a density $\rho$ can only be contained in one subdifferential.
-Otherwise, if $\rho$ is contained in the subdifferential at $v_{1}$ and $v_{2}$, then the function $\Omega^{\beta}$ would not be strictly convex on the line with boundary points $v_{1}$, $v_{2}$.
-By the relation (34) we get that $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$ is single valued.
-This concludes the proof.
-∎
+According to Theorem 7, the functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$ and $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is non-empty for all densities $\rho\in\mathscr{X}_{>0}$.
+By Lemma 13 it is then enough to show that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued.
+Let us assume that there is a $\rho\in\mathscr{X}_{>0}$ for which $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is not single-valued.
+In particular, we assume that $v_{1},v_{2}\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)$ with $v_{1}\neq v_{2}$.
+By Theorem 11 the two Gibbs states $\Gamma_{1}$ and $\Gamma_{2}$ corresponding to $v_{1}$ and $v_{2}$ both yield the density $\rho$.
+Moreover, Proposition 3 states $\Gamma_{1}\neq\Gamma_{2}$.
+We then argue like in the proof of the Hohenberg–Kohn theorem [21] that
+$\Omega^{\beta}(v_{1})$ $=\Omega^{\beta}_{v_{1}}(\Gamma_{1})=\Omega^{\beta}_{v_{2}}(\Gamma_{1})+\langle\rho,v_{1}-v_{2}\rangle>\Omega^{\beta}_{v_{2}}(\Gamma_{2})+\langle\rho,v_{1}-v_{2}\rangle=\Omega^{\beta}(v_{2})+\langle\rho,v_{1}-v_{2}\rangle.$ (39)
+Changing the roles of $v_{1}$ and $v_{2}$ and adding the two inequalities gives the contradiction
+$\Omega^{\beta}(v_{1})+\Omega^{\beta}(v_{2})<\Omega^{\beta}(v_{2})+\Omega^{\beta}(v_{1}).$ (40)
+Hence, $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ must be single-valued. ∎
 VIII Summary and outlook
@@
 This implies that $\rho$ is a minimizer for the variational principle in Eq.
-(15a), where the negative potential is in the subdifferential, i.e., $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$.
+(22a), where the potential is in the negative subdifferential, i.e., $v\in-\partial F^{\beta}_{\mathrm{DM}}(\rho)$.
 The minimizer of Eq.
-(15b), guaranteed to exist by Theorem 11, is then also a minimizer in the Helmholtz functional (Eq.
-(12)).
-This minimizer is unique by Proposition 3 and is equal to the Gibbs state $\Gamma_{v}$.
-We conclude, that $\rho$ is the density of $\Gamma_{v}$ and hence $v$-representable.
+(22b), guaranteed to exist by Theorem 11, is equal to the Gibbs state $\Gamma_{v}$.
+We conclude, that $\rho$ is $v$-representable.
 On the other hand, every density of a Gibbs state $\Gamma_{v}$ with $v\in H^{-1}(\mathbb{T})$ is contained in $\mathscr{I}$ according to Proposition 4 and is strictly positive by Theorem 12.
 In particular, every $v$-representable density is in $\mathscr{X}_{>0}$ and $F^{\beta}_{\mathrm{DM}}$ is not differentiable at densities in $\mathscr{I}\setminus\mathscr{X}_{>0}$.
-On the other hand, Theorem 18 shows that the universal functional $F^{\beta}_{\mathrm{DM}}$ is Gâteaux differentiable at any $\rho\in\mathscr{X}_{>0}$.
+On the other hand, Theorem 14 shows that the thermal universal functional $F^{\beta}_{\mathrm{DM}}$ is Gâteaux differentiable at any $\rho\in\mathscr{X}_{>0}$.
 ∎
-In this work we provided an explicit characterization of the set of uniquely $v$-representable densities on the one-dimensional torus at elevated temperatures and established (Gâteaux) differentiability of the universal functional.
-In the same manner the theory can be developed for any other one-dimensional bounded domain.
+In this work we provided an explicit characterization of the set of uniquely $v$-representable densities on the one-dimensional torus at elevated temperatures and established (Gâteaux) differentiability of the thermal universal functional.
+In the same manner the theory can be developed for any other interval with Neumann boundary conditions.
 As an extensive parameter, the volume needs to be finite, and thus a periodic setting is the natural choice for elevated temperatures.
-The restriction to one dimension remains as the main limitation of the current formulation.
-Another interesting observation is that we do need distributional potentials.
+The restriction to one dimension remains as the main limitation of the current formulation even though some of the results like the uniqueness of the Gibbs state, strict positivity of the density and the uniqueness of the potential also hold in higher dimensions.
+Note that we cannot conclude Gâteaux differentiability in higher dimensions since continuity of the thermal universal functional might not be guaranteed.
+Another interesting observation is that we need to include distributional potentials for achieving a $v$-representability result already in one dimension.
 In DFT often the potential set $L^{3/2}+L^{\infty}$ is considered [28].
-However, for zero temperature and one particle it was already observed by Englisch and Englisch [26] and Chayes et al.
-[12] that relatively nice densities need distributional potentials.
+However, for zero temperature and single particles it was already observed by Englisch and Englisch [11] and Chayes et al.
+[10] that certain densities require distributional potentials for their realization.
 Sutter et al.
-[7] showed that distributional potentials are sufficient for $v$-representability of the set $\mathscr{X}_{>0}$.
-For elevated temperature we further show that distributional potential are also necessary.
-This leads to the clear conjecture that also for higher dimensional domains some sort of distributional potentials are needed.
-However, the continuous Sobolev embedding $H^{1}\hookrightarrow L^{\infty}$ is crucial and it holds only for one-dimensional domains.
-For higher dimension we either need to require higher regularity or higher integrability conditions.
-This means that for the Sobolev space $W^{m,p}$, we need to have $m$ and $p$ large enough.
-The problem is then that not all Gibbs state densities might be contained in $W^{m,p}$ and that a Hamiltonian with a potential in $W^{-m,q}=(W^{m,p})^{\ast}$ might not be self-adjoint.
+[38] showed that distributional potentials of the form Eq.
+(3b) are sufficient for $v$-representability of the set $\mathscr{X}_{>0}$.
+For elevated temperature we further show that such distributional potentials are also necessary.
+Since it cannot be expected that in higher dimensions the space of necessary potentials gets more regular, we conjecture that distributional potentials (probably of a higher class) are also required there if a similar proof of $v$-representability is feasible.
+However, in the current construction, the continuous Sobolev embedding $H^{1}\hookrightarrow L^{\infty}$ is crucial and it holds in one dimension.
+For higher dimensions, we either need to require higher regularity or higher integrability conditions.
+This means that for densities we consider the Sobolev space $W^{m,p}$ and need to have $m$ and/or $p$ large enough.
+The problem is then that not all Gibbs state densities might be contained in $W^{m,p}$ and that a Hamiltonian with a potential in $W^{-m,q}=(W^{m,p})^{\ast}$ will usually not be self-adjoint [20].
 Another open question is how to extend the results to the grand canonical ensemble of variable particle number.
-In this case, the particle-number operator is not bounded anymore and this is currently needed in the proof of Theorem 11 where we show that the weak-* limit yields the same density as the sequence.
+In this case, the particle-number operator is not bounded anymore and this is currently needed in the proof of Theorem 11, where we show that the weak-* limit yields the same density as the sequence.
 Additionally, we do not know if the square root of the Gibbs state density is in $H^{1}$.
-However, we think that the rest of the theory does not fundamentally change and that we get the same conclusion as for the canonical ensemble.
+However, we think that the rest of the theory does not fundamentally change for the grand canonical ensemble and that we retrieve the same conclusions as for the canonical ensemble.
 Acknowledgements.
+We are grateful to Thiago Carvalho Corso for helpful feedback on a previous version of this article.
 MP acknowledges support from the German Research Foundation under Grant SCHI 1476/1-1 and from ERC-2021-STG grant agreement No.
@@
 We still need to show that Proposition 3 holds.
-The result that the unique minimizer of the Helmholtz functional is given by the Gibbs state $\Gamma_{v}=\mathrm{e}^{\beta H_{v}}/Z(v)$ and the relation $\Omega^{\beta}_{v}(\Gamma_{v})=-\beta\log Z(v)$ is shown in standard literature.
-Yet, we give here a different approach than Mermin [11] and a more elaborated version than our previous works [16; 15].
+The result that the unique minimizer of the Helmholtz free energy is given by the Gibbs state $\Gamma_{v}=\mathrm{e}^{\beta H_{v}}/Z(v)$ and the relation $\Omega^{\beta}_{v}(\Gamma_{v})=-\beta\log Z(v)$ is shown in standard literature.
+We give here a more elaborated version than our previous works [17; 37] in the form of a formal proof without going into mathematical subtleties.
+A classical proof can be found in Mermin [30, Appendix] for the grand canonical ensemble but works analogously for the canonical one.
 Proof of Proposition 3.
-To determine the minimizer of the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)=\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$ we make it stationary with respect to small perturbations $\delta\Gamma$ that obey $\operatorname{Tr}\{\delta\Gamma\}=0$.
+To determine the minimizer of the Helmholtz free energy $\Omega^{\beta}_{v}(\Gamma)=\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$ we make it stationary with respect to small perturbations $\delta\Gamma$ that obey $\operatorname{Tr}\{\delta\Gamma\}=0$.
 We compute
@@
 We denote this partition function by $Z_{0}(0)$.
-The eigenstates of the kinetic-energy operator $\frac{1}{2}\Delta$ on $\mathbb{T}$ are Slater determinants built from the functions $\mathrm{e}^{\mathrm{i}2\pi p\cdot x}$ with $p\in\mathbb{Z}$ which have eigenvalues $2\pi^{2}p^{2}$.
+The eigenstates of the kinetic-energy operator $T=-\frac{1}{2}\Delta$ on $\mathbb{T}$ are Slater determinants built from the functions $\mathrm{e}^{\mathrm{i}2\pi p\cdot x}$ with $p\in\mathbb{Z}$ which have eigenvalues $2\pi^{2}p^{2}$.
 We estimate
@@
 To show that $Z(v)$ is finite for a general interaction and a general potential we first show that $\Omega^{\beta}(v)$ is finite.
-For that recall that we assume that the interaction satisfies the KLMN condition (3).
+For that recall that we assume that the interaction satisfies the KLMN condition from Eq.
+(10).
 Moreover, Sutter et al.
-[7, Lem.
-16] showed that any potential $v\in H^{-1}$ fulfills also the KLMN condition $|V(\Psi,\Psi)|\leq aT(\Psi)+b$ where $a>0$ can be made arbitrary small.
+[38, Lem.
+16] showed that any potential $v\in H^{-1}$ also fulfills the KLMN condition $|V(\Psi,\Psi)|\leq a^{\prime}T(\Psi)+b^{\prime}$ where $a^{\prime}>0$ can be made arbitrary small.
 Here $V(\cdot,\cdot)$ is the quadratic form produced by $v\in H^{-1}$.
-Combing these two KLMN conditions guarantees the existence of $0<a<1$ and $b$ such that $\langle\Psi,H_{v}\Psi\rangle\geq(1-a)T(\Psi)-b$ for all $\Psi\in\mathscr{W}$.
+Combing these two KLMN conditions guarantees the existence of new constants $0\leq a<1$ and $b\geq 0$ such that $\langle\Psi,H_{v}\Psi\rangle\geq(1-a)T(\Psi)-b$ for all $\Psi\in\mathscr{W}$.
 This condition translates to density matrices and we get
-$\Omega^{\beta}(v)=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log(\Gamma))\}\geq(1-a)\inf_{\Gamma}\operatorname{Tr}\{\Gamma(T+\big{(}\beta(1-a)\big{)}^{-1}\log(\Gamma))\}-b=-\beta^{-1}\log\big{(}Z_{0}(0)\big{)}-b>-\infty.$ (39)
-Let $|H_{v}=\sum_{j=1}^{\infty}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of the Hamiltonian $H_{v}$ which exist due to the compact embedding of $Q(H_{v})=\big{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$ into $\big{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$.
+$\Omega^{\beta}(v)=\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log(\Gamma))\}\geq(1-a)\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(T+\big(\beta(1-a)\big)^{-1}\log(\Gamma))\}-b=-\beta^{-1}\log\big(Z_{0}(0)\big)-b>-\infty.$ (45)
+Let $|H_{v}=\sum_{j=1}^{\infty}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of the Hamiltonian $H_{v}$ which exist due to the compact embedding of $Q(H_{v})=\big(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\big)^{\wedge N}$ into $\big(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2}\big)^{\wedge N}$.
 We define a sequence of density matrices by
@@
 $\Omega^{\beta}(v)<\operatorname{Tr}\{\Gamma_{n}(H_{v}+\beta^{-1}\log(\Gamma_{n}))=\frac{1}{Z_{n}[v]}\sum\limits_{j=1}^{n}\mathrm{e}^{-\beta\lambda_{j}}(-\beta^{-1}\log Z_{n}[v])=-\beta^{-1}\log Z_{n}[v].$ (47)
-The left-hand side is bounded from below by Eq. (39) and we get that $Z_{n}[v]$ is uniformly bounded from above. This implies $Z(v)<\infty$. ∎
+The left-hand side is bounded from below by Eq. (45) and we get that $Z_{n}[v]$ is uniformly bounded from above. This implies $Z(v)<\infty$.
+Finally, to show that $\Gamma_{v}\in\mathscr{D}$ we need to show $S(\Gamma_{v})<\infty$. We use again $\log\Gamma_{v}=-\beta H_{v}-\log Z(v)$ and show with Lemma 16 that
+$S(\Gamma_{v})=-\operatorname{Tr}\{\Gamma_{v}\log\Gamma_{v}\}=\beta\operatorname{Tr}\{\Gamma_{v}H_{v}\}+\log Z(v)\operatorname{Tr}\{\Gamma_{v}\}<\infty.$ (48)
 Appendix B Proof of Proposition 4
@@
 To do so, note that as explained in the proof of Proposition 3 we have for each potential $v$ two constants $0\leq a<1$ and $b\geq 0$ such that
-$(1-a)T-b\leq H_{v}\leq(1+a)T+b$ (42)
-where $T=-1/2\Delta$ is the kinetic energy operator.
-Moreover, the spectrum of the Hamiltonian $H_{v}$ (and also of $T$) contains only discrete values due to the compact embedding $Q(H_{v})=H^{1}(\mathbb{T}^{N})\hookrightarrow L^{2}(\mathbb{T}^{N})$.
-Then we can express $H_{v}$ and $T$ in spectral decomposition
+$(1-a)T-b\leq H_{v}\leq(1+a)T+b,$ (49)
+where $T=-\frac{1}{2}\Delta$ is the kinetic energy operator.
+Moreover, the Hamiltonian $H_{v}$ (and also $\Delta/2$) has discrete spectrum due to the compact embedding $Q(H_{v})=\bigl(H_{\mathbb{C}}^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}\hookrightarrow\bigl(L^{2}_{\mathbb{C}}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr)^{\wedge N}$.
+We can express $H_{v}$ and $T$ in their respective spectral decomposition
 […]
-where we additionally require $\lambda_{1}\leq\lambda_{2}\leq\ldots$ and $\mu_{1}\leq\mu_{2}\leq\ldots$. Then, we get the following lemma.
+where we additionally require $\lambda_{1}\leq\lambda_{2}\leq\ldots$ and $\mu_{1}\leq\mu_{2}\leq\ldots$. Then we get the following lemma.
 Lemma 15.
-The $n$’th eigenvalue of $H_{v}$ has the lower and upper bound, $(1-a)\mu_{n}-b\leq\lambda_{n}\leq(1+a)\mu_{n}+b$.
+The $n$’th eigenvalue of $H_{v}$ has the following lower and upper bound, $(1-a)\mu_{n}-b\leq\lambda_{n}\leq(1+a)\mu_{n}+b$.
 Proof.
-Let $\phi_{j}$ be the states introduced in Eq. (43). By the min-max-principle for self-adjoint operators and by Eq. (42) we get
+Let $\phi_{j}$ be the states introduced in Eq. (50). By the min-max principle for self-adjoint operators and by Eq. (49) we get
 […]
 […]
-For the lower bound we proceed in the same way but we use the max-min-principle instead to get
+For the lower bound we proceed in the same way but we use the max-min principle instead to get
 […]
 […]
-We use the lower and upper bound of the eigenvalues to proof that the kinetic energy expectation is finite.
+This gives the desired bounds. ∎
+We use the lower and upper bound of the eigenvalues to prove that expectation value of the kinetic energy operator is finite.
 Lemma 16.
-For all, $v\in H^{-1}$ we have $\mathrm{Tr}\{T\Gamma_{v}\}<\infty$.
+For all $v\in H^{-1}(\mathbb{T})$ we have $0\leq\operatorname{Tr}\{T\Gamma_{v}\}<\infty$ and $\infty<\operatorname{Tr}\{H_{v}\Gamma_{v}\}<\infty$.
 Proof.
 The Gibbs state is given by $|\Gamma_{v}=\frac{1}{Z(v)}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$.
-We first show that the energy expectation is finite by using Lemma 19.
+We first show that the expectation value is finite by using Lemma 15.
 That is,
--$\infty$ $<\mathrm{Tr}\{H_{v}\Gamma_{v}\}=\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\lambda_{j}\mathrm{e}^{-\beta\lambda_{j}}\leq\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\big{[}(1+a)\mu_{j}+b\big{]}\mathrm{e}^{-\beta[(1-a)\mu_{j}-b]}$
-$=\frac{(1+a)}{Z(v)}\mathrm{e}^{\beta b}\sum\limits_{j=1}^{\infty}\mu_{j}\mathrm{e}^{-\beta(1-a)\mu_{j}}+\frac{\mathrm{e}^{\beta b}b}{Z(v)}\sum\limits_{j=1}^{\infty}\mathrm{e}^{-\beta(1-a)\mu_{j}}.$
-A similar computation as in Eq. (38) shows that the two sums are finite and we get $\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty$. Using Eq. (42) and positivity of the kinetic energy operator gives
-$-b\leq(1-a)\mathrm{Tr}\{T\Gamma_{v}\}-b\leq\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty,$
+-$\infty$ $<\operatorname{Tr}\{H_{v}\Gamma_{v}\}=\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\lambda_{j}\mathrm{e}^{-\beta\lambda_{j}}\leq\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\big[(1+a)\mu_{j}+b\big]\mathrm{e}^{-\beta[(1-a)\mu_{j}-b]}$ (53)
+$=\frac{\mathrm{e}^{\beta b}(1+a)}{Z(v)}\sum\limits_{j=1}^{\infty}\mu_{j}\mathrm{e}^{-\beta(1-a)\mu_{j}}+\frac{\mathrm{e}^{\beta b}b}{Z(v)}\sum\limits_{j=1}^{\infty}\mathrm{e}^{-\beta(1-a)\mu_{j}}.$
+A similar computation as in Eq. (44) shows that the two sums are finite and we get $\operatorname{Tr}\{H_{v}\Gamma_{v}\}<\infty$. Using Eq. (49) and positivity of the kinetic energy operator gives
+$-b\leq(1-a)\operatorname{Tr}\{T\Gamma_{v}\}-b\leq\operatorname{Tr}\{H_{v}\Gamma_{v}\}<\infty,$ (54)
 which concludes the proof. ∎
@@
 Clearly, $\rho^{n}\to\rho$ in $L^{1}(\mathbb{T})$ and $\|\sqrt{\rho^{n}}\rVert_{2}^{2}\leq N$.
-Since $\int(\nabla\sqrt{\rho})^{2}\mathrm{d}x$ is convex on $\mathscr{I}$ [6, remark after Eq.
-(1.10)], we find by Jensen’s inequality and by Theorem 2
-$\int(\nabla\sqrt{\rho^{n}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}\int(\nabla\sqrt{\rho_{j}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}2T(\psi_{j})\leq 2\mathrm{Tr}\{T\Gamma_{v}\}<\infty,$
+Since $\int(\nabla\sqrt{\rho})^{2}\mathrm{d}x$ is convex on $\mathscr{I}$ [28, remark after Eq.
+(1.10)], we find by Jensen’s inequality and by Theorem 2 that
+$\int(\nabla\sqrt{\rho^{n}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}\int(\nabla\sqrt{\rho_{j}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}2T(\psi_{j})\leq 2\operatorname{Tr}\{T\Gamma_{v}\}<\infty,$ (55)
 where we get a uniform bound by Lemma 16.
-In particular, $\sqrt{\rho^{n}}$ is a bounded sequence in $H^{1}(\mathbb{T})$ and by the Banach–Alaoglu theorem, there is a subsequence (also denoted by $\sqrt{\rho^{n}}$) and $f\in H^{1}(\mathbb{T})$ such that $\sqrt{\rho^{n}}\rightharpoonup f$ weakly in $H^{1}$.
+In particular, $\sqrt{\rho^{n}}$ is a bounded sequence in $H^{1}(\mathbb{T})$ and by the Banach–Alaoglu theorem there is a subsequence (also denoted by $\sqrt{\rho^{n}}$) and $f\in H^{1}(\mathbb{T})$ such that $\sqrt{\rho^{n}}\rightharpoonup f$ weakly in $H^{1}(\mathbb{T})$.
 But since $H^{1}(\mathbb{T})$ is compactly embedded in $L^{2}(\mathbb{T})$ there is also a subsequence (again denoted by $\sqrt{\rho^{n}}$) such that $\sqrt{\rho^{n}}\to f$ strongly in $L^{2}(\mathbb{T})$.
 Then, we have
-$\|\rho^{n}-f^{2}\rVert_{1}=\int(\sqrt{\rho^{n}}-f)(\sqrt{\rho^{n}}+f)\mathrm{d}x\leq\|\sqrt{\rho^{n}}-f\rVert_{2}\|\sqrt{\rho^{n}}+f\rVert_{2}\to 0$
-since $\|\sqrt{\rho^{n}}+f\rVert_{2}$ is bounded. This means that $\rho^{n}\to f^{2}$ strongly in $L^{1}$ and we conclude $\sqrt{\rho}=f\in H^{1}(\mathbb{T})$. ∎
+$\|\rho^{n}-f^{2}\rVert_{1}=\int(\sqrt{\rho^{n}}-f)(\sqrt{\rho^{n}}+f)\mathrm{d}x\leq\|\sqrt{\rho^{n}}-f\rVert_{2}\|\sqrt{\rho^{n}}+f\rVert_{2}\to 0$ (56)
+since $\|\sqrt{\rho^{n}}+f\rVert_{2}$ is bounded. This means that $\rho^{n}\to f^{2}$ strongly in $L^{1}(\mathbb{T})$ and we conclude $\sqrt{\rho}=f\in H^{1}(\mathbb{T})$. ∎